C I T A T I O N S

of the scientific publications^{∗} of Bozhidar Zakhariev ILIEV ^{†} ^{‡} ^{�}

The work [1] is cited by:

1. (1) S. Manoff, Deviation equations of Synge and Schild over spaces with affine connections and metrics, Intern. J. Mod. Phys. A, vol. 16, No. 6, pp. 1109-1122, 2001; http://arXiv.org e-Print archive, E-print No. gr-qc/0012015, 2000.

2. (2) Sawa Manoff, Geometry and mechanics in different models of space-time: Geometry and kinematics, Nova science Publishers, Inc., Yew York, 2002

3. (3) Sawa Manoff, Geometry and mechanics in different models of space-time: Dynamics and applications, Nova science Publishers, Inc., Yew York, 2002

The work [2] is cited by:

1. (4) S. Manoff, Deviation equations of Synge and Schild over spaces with affine connections and metrics, Intern. J. Mod. Phys. A, vol. 16, No. 6, pp. 1109-1122, 2001; http://arXiv.org e-Print archive, E-print No. gr-qc/0012015, 2000.

2. (5) Sawa Manoff, Geometry and mechanics in different models of space-time: Geometry and kinematics, Nova science Publishers, Inc., Yew York, 2002

3. (6) Sawa Manoff, Geometry and mechanics in different models of space-time: Dynamics and applications, Nova science Publishers, Inc., Yew York, 2002

4. (7) S. Manoff, Mechanics of Continuous Media in (L̄_{n}, g)-Spaces. III. Relative Accelerations, http://arXiv.org e-Print archive, E-print No. gr-qc/0204003, 2002.

The work [3] is cited by:

1. (8) Christian Gebken, Conformal Geometric Algebra in Stochastic Optimization Problems of 3D-Vision Applications, Ph.D. Thesis, der Technischen Fakultat der Christian-Albrechts-Universit�at zu Kiel, Kiel, 2010.

The work [3] is cited by:

1. (9) Christian Gebken, Conformal Geometric Algebra in Stochastic Optimization Problems of 3D-Vision Applications, Ph.D. Thesis, der Technischen Fakultat der Christian-Albrechts-Universit�at zu Kiel, Kiel, 2010.

The work [4] is cited by:

1. (10) S. Manoff, Kinematics of vector fields, in: “Complex Structure and Vector Fields”, World Scientific Publ., Singapore, 1995, pp. 61-113.

2. (11) S. Manoff, Geodesic and autoparallel equation over differentiable manifolds, Int. J. Modern Phys. A, vol. 11, No. 21, 1996, pp. 3849-3874.

∗The list of the cited publications is at the end of this document, p. 28.

†Laboratory of Mathematical Modeling in Physics, Institute for Nuclear Research and Nuclear Energy, Bulgarian Academy of Sciences, Boul. Tzarigradsko chaussée 72, 1784 Sofia, Bulgaria, tel.:+359-2-979-5646, fax:+359-2-9753619.

‡E-mail address: bozho@inrne.bas.bg �URL: http://theo.inrne.bas.bg/∼bozho/

\underline{List of citations of the scientific publications of Bozhidar Z. Iliev} 2

3. (12) F. Gronwald, F. W. Hehl, On the gauge aspects of gravity, Proc. of the 14th Course of the School of Cosmology and Gravitation on Quantum Gravity, Erice, Italy, May 1995, P.G. Bergmann, V. de Sabbata, and H.-J. Treder, eds., World Scientific, Singapore, 1996.

4. (13) S. Manoff, Lagrangian formalism for tensor fields, in: Topics in complex analysis, differential geometry and mathematical physics, eds. K. Sekigawa and S. Dimiev, World Scientific Pbl., 1997, pp. 177-218.

5. (14) S. Manoff, Fermi-Walker transport over spaces with affine connection and metric, JINR Rapid Communications No. 1(81)-97, Dubna, 1997.

6. (15) S. Manoff, Fermi derivative and Fermi-Walker transports over (\overline{L}_{n}, g)-spaces, Int. J. Modern Phys. A, vol. 13, No. 25, 1998, pp. 4289-4308.

7. (16) S. Manoff, A. Kolarov, B. Dimitrov, (\overline{L}_{n}, g)-spaces. General relativity over \overline{V}_{4}- spaces, Communication JINR, E5-98-184, Dubna, 1998.

8. (17) S. Manoff, Lagrangian theory of tensor fields over spaces with contravariant and covariant affine connections and metrics and its applications to Einstein’s theory of gravitation in \overline{L}_{4} spaces, Acta applicandae mathematicae, vol. 55, pp. 51–125, 1999.

9. (18) S. Manoff, Frames of reference in spaces with affine connections and metrics, Classical and quantum gravity, vol. 18, No. 6, 2001, pp. 1111–1125. http://arXiv.org e-Print archive, E-print No. gr-qc/9908061, 1999.

10. (19) S. Manoff, Auto-parallel equation as Euler-Lagrange’s equation in spaces with affine connections and metrics, Gen. Rel. Grav., vol. 32, No. 8, pp. 1559–1582, 2000. http://arXiv.org e-Print archive, E-print No. gr-qc/0010048, 2000.

11. (20) S. Manoff, B. Dimitrov, Weyl’s spaces with shear-free and expansion-free conformal Killing vectors and the motion of a free spinless test particle, http://arXiv.org e-Print archive, E-print No. gr-qc/0011045, 2000.

12. (21) S. Manoff, Lagrangian formalism for tensor fields, http://arXiv.org e-Print archive, E-print No. gr-qc/0007056, 2000.

13. (22) S. Manoff, Mechanics of continuous media in (L̄_{n}, g) spaces. 1. Introduction and mathematical tools. http://arXiv.org e-Print archive, E-print No. gr-qc/0203016, March 2002.

14. (23) Sawa Manoff, Geometry and mechanics in different models of space-time: Geometry and kinematics, Nova science Publishers, Inc., Yew York, 2002

15. (24) Sawa Manoff, Geometry and mechanics in different models of space-time: Dynamics and applications, Nova science Publishers, Inc., Yew York, 2002

16. (25) S. Manoff, B. Dimitrov, Flows and particles with shear - free and expansion - free velocities in (L̄_{n}, g) and weyl’s spaces, Classical and quantum gravity, vol. 19, pp. 4377-4398,2002 http://arXiv.org e-Print archive, E-print No. gr-qc/0207060, July 2002

17. (26) R. Aldrovandi, P. B. Barros, J. G. Pereira, The equivalence principle revisited, Foundations of physics, vol. 33, No. 4, pp. 545–575, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0212034, 2002.33(4): 545–575; Apr 2003 18. (27) Arminjon, M., Reifler, F, General reference frames and their associated space

manifolds, International Journal of Geometric Methods in Modern Physics vol. 8, No. 1, pp. 155–165, 2011, DOI: 10.1142/S0219887811005051.

19. (28) Gagik Ter Kazarian, Probing the origin of inertia behind spacetime deformation, http://arXiv.org e-Print archive, E-print No. arXiv:1105.5932 [gr-qc], May 2011.

20. (29) Gagik Ter Kazarian, Spacetime Deformation-Induced Inertia Effects, Advances in Mathematical Physics (Hindawi Publishing Corporation), Volume 2012, Article ID 692030, 41 pages, 2012; doi:10.1155/2012/692030.

The work [5] is cited by:

1. (30) S. Manoff, Kinematics of vector fields, in: “Complex Structure and Vector Fields”, World Scientific Publ., Singapore, 1995, pp. 61-113.

2. (31) S. Manoff, Geodesic and autoparallel equation over differentiable manifolds, Int. J. Modern Phys. A, vol. 11, No. 21, 1996, pp. 3849-3874.

3. (32) F. Gronwald, F. W. Hehl, On the gauge aspects of gravity, Proc. of the 14th Course of the School of Cosmology and Gravitation on Quantum Gravity, Erice, Italy, May 1995, P.G. Bergmann, V. de Sabbata, and H.-J. Treder, eds., World Scientific, Singapore, 1996.

4. (33) S.Manoff, Lagrangian formalism for tensor fields, in: Topics in complex analysis, differential geometry and mathematical physics, eds. K. Sekigawa and S. Dimiev, World Scientific Pbl., 1997, pp. 177-218.

5. (34) S. Manoff, Fermi-Walker transport over spaces with affine connection and metric, JINR Rapid Communications No. 1(81)-97, Dubna, 1997.

6. (35) S. Manoff, Fermi derivative and Fermi-Walker transports over (\overline{L}_{n}, g)-spaces, Int. J. Modern Phys. A, vol. 13, No. 25, 1998, pp. 4289-4308.

7. (36) S. Manoff, A. Kolarov, B. Dimitrov, (\overline{L}_{n}, g)-spaces. General relativity over \overline{V}_{4}- spaces, Communication JINR, E5-98-184, Dubna, 1998.

8. (37) S. Manoff, Lagrangian theory of tensor fields over spaces with contravariant and covariant affine connections and metrics and its applications to Einstein’s theory of gravitation in \overline{L}_{4} spaces, Acta applicandae mathematicae, vol. 55, pp. 51–125, 1999.

9. (38) S. Manoff, Frames of reference in spaces with affine connections and metrics, Classical and quantum gravity, vol. 18, No. 6, 2001, pp. 1111–1126. http://arXiv.org e-Print archive, E-print No. gr-qc/9908061, 1999.

10. (39) S. Manoff, Auto-parallel equation as Euler-Lagrange’s equation in spaces with affine connections and metrics, Gen. Rel. Grav., vol. 32, No. 8, pp. 1559–1582, 2000. http://arXiv.org e-Print archive, E-print No. gr-qc/0010048, 2000.

11. (40) S. Manoff, B. Dimitrov, Weyl’s spaces with shear-free and expansion-free conformal Killing vectors and the motion of a free spinless test particle, http://arXiv.org e-Print archive, E-print No. gr-qc/0011045, 2000.

12. (41) S. Manoff, Lagrangian formalism for tensor fields, http://arXiv.org e-Print archive, E-print No. gr-qc/0007056, 2000.

13. (42) S. Manoff, B. Dimitrov, Flows and particles with shear - free and expansion - free velocities in (L̄_{n}, g) and weyl’s spaces, Classical and quantum gravity, vol. 19, pp. 4377-4398,2002 http://arXiv.org e-Print archive, E-print No. gr-qc/0207060, July 2002

14. (43) S. Manoff, Mechanics of continuous media in (L̄_{n}, g) spaces. 1. Introduction and mathematical tools. http://arXiv.org e-Print archive, E-print No. gr-qc/0203016, March 2002.

15. (44) R. Aldrovandi, P. B. Barros, J. G. Pereira, The equivalence principle revisited, Foundations of physics, vol. 33, No. 4, pp. 545–575, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0212034, 2002.

16. (45) Sawa Manoff, Geometry and mechanics in different models of space-time: Geometry and kinematics, Nova science Publishers, Inc., Yew York, 2002

17. (46) Sawa Manoff, Geometry and mechanics in different models of space-time: Dynamics and applications, Nova science Publishers, Inc., Yew York, 2002

18. (47) Gagik Ter Kazarian, Probing the origin of inertia behind spacetime deformation, http://arXiv.org e-Print archive, E-print No. arXiv:1105.5932 [gr-qc], May 2011.

The work [6] is cited by:

1. (48) S. Manoff, Kinematics of vector fields, in: “Complex Structure and Vector Fields”, World Scientific Publ., Singapore, 1995, pp. 61-113.

2. (49) S. Manoff, Geodesic and autoparallel equation over differentiable manifolds, Int. J. Modern Phys. A, vol. 11, No. 21, 1996, pp. 3849-3874.

3. (50) F. Gronwald, F. W. Hehl, On the gauge aspects of gravity, Proc. of the 14th Course of the School of Cosmology and Gravitation on Quantum Gravity, Erice, Italy, May 1995, P.G. Bergmann, V. de Sabbata, and H.-J. Treder, eds., World Scientific, Singapore, 1996.

4. (51) S.Manoff, Lagrangian formalism for tensor fields, in: Topics in complex analysis, differential geometry and mathematical physics, eds. K. Sekigawa and S. Dimiev, World Scientific Pbl., 1997, pp. 177-218.

5. (52) S. Manoff, Fermi-Walker transport over spaces with affine connection and metric, JINR Rapid Communications No. 1(81)-97, Dubna, 1997.

6. (53) S. Manoff, Fermi derivative and Fermi-Walker transports over (\overline{L}_{n}, g)-spaces, Int. J. Modern Phys. A, vol. 13, No. 25, 1998, pp. 4289-4308.

7. (54) S. Manoff, A. Kolarov, B. Dimitrov, (\overline{L}_{n}, g)-spaces. General relativity over \overline{V}_{4}- spaces, Communication JINR, E5-98-184, Dubna, 1998.

8. (55) S. Manoff, Lagrangian theory of tensor fields over spaces with contravariant and covariant affine connections and metrics and its applications to Einstein’s theory of gravitation in \overline{L}_{4} spaces, Acta applicandae mathematicae, vol. 55, pp. 51–125, 1999.

9. (56) S. Manoff, Frames of reference in spaces with affine connections and metrics, Classical and quantum gravity, vol. 18, No. 6, 2001, pp. 1111–1126. http://arXiv.org e-Print archive, E-print No. gr-qc/9908061, 1999.

10. (57) S. Manoff, Auto-parallel equation as Euler-Lagrange’s equation in spaces with affine connections and metrics, Gen. Rel. Grav., vol. 32, No. 8, pp. 1559–1582, 2000. http://arXiv.org e-Print archive, E-print No. gr-qc/0010048, 2000.

11. (58) S. Manoff, B. Dimitrov, Weyl’s spaces with shear-free and expansion-free conformal Killing vectors and the motion of a free spinless test particle, http://arXiv.org e-Print archive, E-print No. gr-qc/0011045, 2000.

12. (59) S. Manoff, Lagrangian formalism for tensor fields, http://arXiv.org e-Print archive, E-print No. gr-qc/0007056, 2000.

13. (60) S. Manoff, Mechanics of continuous media in (L̄_{n}, g) spaces. 1. Introduction and mathematical tools. http://arXiv.org e-Print archive, E-print No. gr-qc/0203016, March 2002.

14. (61) R. Aldrovandi, P. B. Barros, J. G. Pereira, The equivalence principle revisited, Foundations of physics, vol. 33, No. 4, pp. 545–575, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0212034, 2002.

15. (62) Sawa Manoff, Geometry and mechanics in different models of space-time: Geometry and kinematics, Nova science Publishers, Inc., Yew York, 2002

16. (63) Sawa Manoff, Geometry and mechanics in different models of space-time: Dynamics and applications, Nova science Publishers, Inc., Yew York, 2002

17. (64) Andrew J.S. Hamilton and Pedro P. Avelino, The Physics of the relativistic counterstreaming instability that drives mass inflation inside black holes, Phys.Rept. 495, pp. 1-32, 2010, http://arXiv.org e-Print archive, E-print No. arXiv:0811.1926 [gr-qc], 2010.

18. (65) Gagik Ter Kazarian, Probing the origin of inertia behind spacetime deformation, http://arXiv.org e-Print archive, E-print No. arXiv:1105.5932 [gr-qc], May 2011.

The work [7] is cited by:

1. (66) D. D. Ferrante Geometric calculus and the fibre bundle description of quantum mechanics, Preprint BROWN-HET-1307, March 2002; http://arXiv.org e-Print archive, E-print No. math-ph/0204024, 2002.

The work [8] is cited by:

1. (67) D. D. Ferrante Geometric calculus and the fibre bundle description of quantum mechanics, Preprint BROWN-HET-1307, March 2002; http://arXiv.org e-Print archive, E-print No. math-ph/0204024, 2002.

The work [9] is cited by:

1. (68) D. D. Ferrante Geometric calculus and the fibre bundle description of quantum mechanics, Preprint BROWN-HET-1307, March 2002; http://arXiv.org e-Print archive, E-print No. math-ph/0204024, 2002.

The work [10] is cited by:

1. (69) D. D. Ferrante Geometric calculus and the fibre bundle description of quantum mechanics, Preprint BROWN-HET-1307, March 2002; http://arXiv.org e-Print archive, E-print No. math-ph/0204024, 2002.

2. (70) Daniel D. Ferrante, Generalized Bundle Quantum Mechanics, Preprint BROWN-HET-1374, 2003; http://arXiv.org e-Print archive, E-print No. math-ph/0308004, August 2003.

The work [11] is cited by:

1. (71) D. D. Ferrante Geometric calculus and the fibre bundle description of quantum mechanics, Preprint BROWN-HET-1307, March 2002; http://arXiv.org e-Print archive, E-print No. math-ph/0204024, 2002.

The work [12] is cited by:

1. (72) D. D. Ferrante Geometric calculus and the fibre bundle description of quantum mechanics, Preprint BROWN-HET-1307, March 2002; http://arXiv.org e-Print archive, E-print No. math-ph/0204024, 2002.

2. (73) Daniel D. Ferrante, Generalized Bundle Quantum Mechanics, Preprint BROWN-HET-1374, 2003; http://arXiv.org e-Print archive, E-print No. math-ph/0308004, August 2003.

The work [13] is cited by:

1. (74) D. D. Ferrante Geometric calculus and the fibre bundle description of quantum mechanics, Preprint BROWN-HET-1307, March 2002; http://arXiv.org e-Print archive, E-print No. math-ph/0204024, 2002.

2. (75) Daniel D. Ferrante, Generalized Bundle Quantum Mechanics, Preprint BROWN-HET-1374, 2003; http://arXiv.org e-Print archive, E-print No. math-ph/0308004, August 2003.

The work [14] is cited by:

1. (76) Dirk Puetzfeld and Yuri N. Obukhov, Deviation equation in Riemann-Cartan spacetime, http://arXiv.org e-Print archive, E-print No. arXiv:1804.11106 [gr-qc], April 2018.

2. (77) Yuri N. Obukhov and Dirk Puetzfeld, Measuring the gravitational field in General Relativity: From deviation equations and the gravitational compass to relativistic clock gradiometry, http://arXiv.org e-Print archive, E-print No. arXiv:1811.10474 [gr-qc], 21 November 2018

3. (78) Yuri N. Obukhov and Dirk Puetzfeld, Measuring the Gravitational Field in General Relativity: From Deviation Equations and the Gravitational Compass to Relativistic Clock Gradiometry,

Chapter in: Puetzfeld D. and Lдmmerzahl C. (eds) Relativistic Geodesy. Fundamental Theories of Physics, vol 196. Springer, Cham, pp. 87–130, 10 February 2019 DOI: 10.1007/978-3-030-11500-5_3, Print ISBN 978-3-030-11499-2, Online ISBN 978-3-030-11500-5.

The work [15] is cited by:

1. (79) Sérgio Costa Ulhoa, O momento angular do campo gravitacional e o grupo de Poincaré, 2009. 96 f. Tese (Doutorado em Fнsica)-Universidade de Brasнlia, Brasнlia, 2009.

The work [16] is cited by:

1. (80) Friedrich W. Hehl, Alternative gravitational theories in four dimensions, preprint COLOGNE-THP-HE-97-19 (Cologne U.), Dec 1997. 10p. http://arXiv.org e-Print archive, E-print No. gr-qc/9712096, December 1997.

2. (81) Friedrich W. Hehl, Alternative Gravitational Theories in Four Dimensions Report of parallel session chair in: Proc. 8th M. Grossmann Meeting, T. Piran, ed., World Scientific, Singapore 1998. Talk given at the 8th Marcel Grossmann Meeting on Recent Developments in Theoretical and Experimental General Relativity, Gravitation and Relativistic Field Theories (MG 8), Jerusalem, Israel, 22-27 Jun 1997; Recent developments in theoretical and experimental general relativity, gravitation, and relativistic field theories, Pt.A* 423-432. http://arXiv.org e-Print archive, E-print No. gr-qc/9712096, 1997. Report of parallel session chair in: Proc. 8th M. Grossmann Meeting, T. Piran (ed.) World Scientific, Singapore 1998.

3. (82) Uwe Munc, Uber teleparallele Gravitationstheorien, Ph.D. Thesis, Institut fur Theoretische Physik Universit�at zu Koln, 1997.

4. (83) Roman Plyatsko, Gravitational ultrarelativistic spin orbit interaction and the weak equivalence principle, Phys.Rev.D, vol. 5808, p. 4031, 1998.

5. (84) S. Manoff, Fermi derivative and Fermi-Walker transports over (\overline{L}_{n}, g)-spaces, Int. J. Modern Phys. A, vol. 13, No. 25, 1998, pp. 4289-4308.

6. (85) S. Manoff, A. Kolarov, B. Dimitrov, (\overline{L}_{n}, g)-spaces. General relativity over \overline{V}_{4}- spaces, Communication JINR, E5-98-184, Dubna, 1998.

7. (86) H. C. Rosu, Classical and quantum inertia: A matter of principles, Gravitation and Cosmology, vol. 5, No. 2 (18) (June 1999), pp. 81-91; http://arXiv.org e-Print archive, E-print No. gr-qc/9412012, 1994.

8. (87) S. Manoff, Relative velocity and relative acceleration induced by the torsion in (pseudo) Riemannian spaces with torsion and in spaces with an affine connection and metrics, http://arXiv.org e-Print archive, E-print No. gr-qc/0005027, 2000.

9. (88) S. Manoff, Spaces with contravariant and covariant affine connections and metrics, Part. Nucl. vol. 30, 517-549, 1999; Fiz. Elem. Chast. Atom. Yadra vol. 30, No. 3 1211-1269, 1999. http://arXiv.org e-Print archive, E-print No. gr-qc/0006024, 2000.

10. (89) S. Manoff, Lagrangian theory of tensor fields over spaces with contravariant and covariant affine connections and metrics and its applications to Einstein’s theory of gravitation in \overline{L}_{4} spaces, Acta applicandae mathematicae, vol. 55, pp. 51–125, 1999.

11. (90) S. Manoff, Auto-parallel equation as Euler-Lagrange’s equation in spaces with affine connections and metrics, Gen. Rel. Grav., vol. 32, No. 8, pp. 1559–1582, 2000. http://arXiv.org e-Print archive, E-print No. gr-qc/0010048, 2000.

12. (91) Friedrich W. Hehl, Yuri N. Obukhov, How does the electromagnetic field couple to gravity, in particular to metric, nonmetricity, torsion, and curvature?, Contribution to "Testing Relativistic Gravity in Space: Gyroscopes, Clocks, Interferometers: Testing General Relativity in Space Proceedings of the 220th Heraeus-Seminar, Bad Honnef, Germany, 22 - 27 August 1999 in Bad Honnef, C. Laemmerzahl et al. (eds.), Springer, Berlin, 2000. http://arXiv.org e-Print archive, E-print No. gr-qc/0001010, 2000. Lect. Notes Phys., vol. 562, p. 479, 2001.

13. (92) S. Manoff, B. Dimitrov, Weyl’s spaces with shear-free and expansion-free conformal Killing vectors and the motion of a free spinless test particle, http://arXiv.org e-Print archive, E-print No. gr-qc/0011045, 2000.

14. (93) S. Manoff, Frames of reference in spaces with affine connections and metrics, Classical and quantum gravity, vol. 18, No. 6, 2001, pp. 1111–1126. http://arXiv.org e-Print archive, E-print No. gr-qc/9908061, 1999.

15. (94) S. Manoff, Deviation equations of Synge and Schild over spaces with affine connections and metrics, Intern. J. Mod. Phys. A, vol. 16, No. 6, pp. 1109-1122, 2001; http://arXiv.org e-Print archive, E-print No. gr-qc/0012015, 2000.

16. (95) S. Manoff, Deviation operator and deviation equations over spaces with affine connections and metrics, Journal of Geometry and Physics, Vol. 39, No. 4, October 2001, pp. 337-350.

17. (96) S. Manoff, B. Dimitrov, Flows and particles with shear - free and expansion - free velocities in (L̄_{n}, g) and weyl’s spaces, Classical and quantum gravity, vol. 19, pp. 4377-4398,2002 http://arXiv.org e-Print archive, E-print No. gr-qc/0207060, July 2002

18. (97) S. Manoff, Mechanics of continuous media in (L̄_{n}, g) spaces. 1. Introduction and mathematical tools. http://arXiv.org e-Print archive, E-print No. gr-qc/0203016, March 2002.

19. (98) R. Aldrovandi, P. B. Barros, J. G. Pereira, The equivalence principle revisited, Foundations of physics, vol. 33, No. 4, pp. 545–575, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0212034, 2002.

20. (99) S. Manoff, B. Dimitrov, On the existence of a gyroscope in spaces with affine connections and metrics, General Relativity and Gravitation, vol. 35, pp. 25–33, 2003, http://arXiv.org e-Print archive, E-print No. gr-qc/0012011, 2000.

21. (100) Sawa Manoff, Geometry and mechanics in different models of space-time: Geometry and kinematics, Nova science Publishers, Inc., Yew York, 2002

22. (101) Sawa Manoff, Geometry and mechanics in different models of space-time: Dynamics and applications, Nova science Publishers, Inc., Yew York, 2002

23. (102) R. Aldrovandi, P. B. Barros, J. G. Pereira, Gravitation as anholonomy, General Relativity and Gravitation, vol. 35, No. 6, pp. 991–1005, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0301077, 2003.

24. (103) M. Blagojević, Three lectures on Poincaré gauge theory, Based on lectures presented at II Summer School in Modern Mathematical Physics, Kopaonik, Yugoslavia, 1-12 September 2002., 2003.

25. (104) S. Manoff, Doppler effect and Hubble effect in different models of space-time in the case of auto-parallel motion of the observer, http://arXiv.org e-Print archive, E-print No. gr-qc/0401012, January 2004.

26. (105) S. Manoff, Determination of the velocity of an emitter in spaces with affine connections and metrics, In “Applications and development of Lobachevski ideas in modern physics”, Proceedings of the International seminar dedicated to the 75th birthday of professor Nikolai A. Chernikov, Dubna 25–27 February 2004, Dubna, 2004. http://arXiv.org e-Print archive, E-print No. gr-qc/0403025, March 2004.

27. (106) S. Manoff, Standard emitters (clocks) and calibrated standard emitters (clocks) in spaces with affine connections and metrics, http://arXiv.org e-Print archive, E-print No. gr-qc/0505061, May 2005.

28. (107) R. Aldrovandi, A. L. Barbosa, Connection space approach to ambiguities of gauge theories, International Journal of Mathematics and Mathematical Sciences, Vol. 2005, No. 15, pp. 2365–2371, 2005.

29. (108) Mayeul Arminjon, Dirac-type equations in a gravitational field, with vector wave function, Found. Phys., Vol. 38, pp. 1020-1045, 2008, DOI: 10.1007/s10701-008-9249-6; http://arXiv.org e-Print archive, E-print No. arXiv:gr-qc/0702048, 2007 and 2008.

30. (109) Mayeul Arminjon, Dirac-type equations in a gravitational field, with vector wave function, Foundations Physics, vol. 38, pp. 1020�1045, 2008, DOI 10.1007/s10701-008-9249-6.

31. (110) Pawe� Laskos-Grabowski, The Einstein�Cartan theory: the meaning and consequences of torsion, Master�s Thesis (Advisor: Professor Jerzy Kowalski-Glikman), University of Wroc�aw, Faculty of Physics and Astronomy. Theoretical physics, August 11, 2009, http://www.ift.uni.wroc.pl/ plg/download/msc.pdf

32. (111) Mayeul Arminjon, Frank Reifler, General reference frames and their associated space manifolds, http://arXiv.org e-Print archive, E-print No. arXiv:1003.3521 [gr-qc], April 2010.

33. (112) James M. Nester, Normal frames for general connections, Ann. Phys. (Berlin), vol. 19, No. 1�2, pp. 45�52, 2010, DOI: 10.1002/andp.200910373.

34. (113) Socolovsky M., Locally inertial coordinates with totally antisymmetric torsion, http://arXiv.org e-Print archive, E-print No. arXiv:1009.3979v2, v1 September 2010., v2 April 2011.

35. (114) Aldrovandi R. and J. G. Pereira, An Introduction to Teleparallel gravity, Instituto de Fisica Teorica, Universidade Estadual Paulista, Sao Paulo, Brazil, http://www.ift.unesp.br/gcg/tele.pdf, 2010.

36. (115) Arminjon, M., Reifler, F, General reference frames and their associated space manifolds, International Journal of Geometric Methods in Modern Physics vol. 8, No. 1, pp. 155–165, 2011, DOI: 10.1142/S0219887811005051.

37. (116) Giglio J. F. T. and W. A. Rodrigues Jr, Locally Inertial Reference Frames in Lorentzian and Riemann-Cartan Spacetimes, Analen der physik, Vol. 524, Issue 5, pp. 302–310, 2012, DOI: 10.1002/andp.201100302. http://arXiv.org e-Print archive, E-print No. arXiv:1111.2206v2 [math-ph], November 2011.

38. (117) Garecki Janusz, Is torsion needed in a theory of gravity? A reappraisal, http://arXiv.org e-Print archive, E-print No. arXiv:1110.4251v1 [gr-qc], October 2011.

39. (118) Aldrovandi R. and J. G. Pereira, An Introduction to teleparallel gravity, Fundamental Theories of Physics, Vol. 173, pp. 179–186, 2013, DOI 10.1007/978-94-007-5143-9_18.

40. (119) R. Aldrovandi, J. G. Pereira, An Introduction to teleparallel gravity, Instituto de Fisica Teorica, Universidade Estadual Paulista, Sao Paulo, Brazil, 2013.

41. (120) Ralitza K. Kovacheva, Montel’s Type Results and Zero Distribution of Sequences of rational Functions, Bulletin de la Societe des sciences et des letters de Lodz, Se�rie: Recherches sur les de�formations, In memory of Professor Roman Stanislaw Ingarden, Volume LXI, No. 3, pp. 7�-13, Lodz, 2011.

42. (121) Friedrich W. Hehl and Yuri N. Obukhov Conservation of energy-momentum of matter as the basis for the gauge theory of gravitation, In One Hundred Years of Gauge Theory, pp. 217-252, DOI: 10.1007/978-3-030-51197-5_10 In Fundamental Theories of Physics book (FTPH), volume 199, Springer, 2020, http://arXiv.org e-Print archive, E-print No. arXiv:1909.01791 [gr-qc], 1 September 2019.

The work [17] is cited by:

1. (122) D. D. Ferrante Geometric calculus and the fibre bundle description of quantum mechanics, Preprint BROWN-HET-1307, March 2002; http://arXiv.org e-Print archive, E-print No. math-ph/0204024, 2002.

The work [18] is cited by:

1. (123) D. D. Ferrante Geometric calculus and the fibre bundle description of quantum mechanics, Preprint BROWN-HET-1307, March 2002; http://arXiv.org e-Print archive, E-print No. math-ph/0204024, 2002.

2. (124) Daniel D. Ferrante, Generalized Bundle Quantum Mechanics, Preprint BROWN-HET-1374, 2003; http://arXiv.org e-Print archive, E-print No. math-ph/0308004, August 2003.

The work [19] is cited by:

1. (125) Uwe Munc, Uber teleparallele Gravitationstheorien, Ph.D. Thesis, Institut fur Theoretische Physik Universit�at zu Koln, 1997.

2. (126) Roman Plyatsko, Gravitational ultrarelativistic spin orbit interaction and the weak equivalence principle, Phys.Rev.D, vol. 5808, p. 4031, 1998.

3. (127) S. Manoff, Fermi derivative and Fermi-Walker transports over (\overline{L}_{n}, g)-spaces, Int. J. Modern Phys. A, vol. 13, No. 25, 1998, pp. 4289-4308.

4. (128) S. Manoff, A. Kolarov, B. Dimitrov, (\overline{L}_{n}, g)-spaces. General relativity over \overline{V}_{4}- spaces, Communication JINR, E5-98-184, Dubna, 1998.

5. (129) H. C. Rosu, Classical and quantum inertia: A matter of principles, Gravitation and Cosmology, vol. 5, No. 2 (18) (June 1999), pp. 81-91; http://arXiv.org e-Print archive, E-print No. gr-qc/9412012, 1994.

6. (130) S. Manoff, Relative velocity and relative acceleration induced by the torsion in (pseudo) Riemannian spaces with torsion and in spaces with an affine connection and metrics, http://arXiv.org e-Print archive, E-print No. gr-qc/0005027, 2000.

7. (131) S. Manoff, Spaces with contravariant and covariant affine connections and metrics, Part. Nucl. vol. 30, 517-549, 1999; Fiz. Elem. Chast. Atom. Yadra vol. 30, No. 3 1211-1269, 1999. http://arXiv.org e-Print archive, E-print No. gr-qc/0006024, 2000.

8. (132) S. Manoff, Lagrangian theory of tensor fields over spaces with contravariant and covariant affine connections and metrics and its applications to Einstein’s theory of gravitation in \overline{L}_{4} spaces, Acta applicandae mathematicae, vol. 55, pp. 51–125, 1999.

9. (133) S. Manoff, Auto-parallel equation as Euler-Lagrange’s equation in spaces with affine connections and metrics, Gen. Rel. Grav., vol. 32, No. 8, pp. 1559–1582, 2000, http://arXiv.org e-Print archive, E-print No. gr-qc/0010048, 2000.

10. (134) S. Manoff, B. Dimitrov, Weyl’s spaces with shear-free and expansion-free conformal Killing vectors and the motion of a free spinless test particle, http://arXiv.org e-Print archive, E-print No. gr-qc/0011045, 2000.

11. (135) S. Manoff, Frames of reference in spaces with affine connections and metrics, Classical and quantum gravity, vol. 18, No. 6, 2001, pp. 1111–1126. http://arXiv.org e-Print archive, E-print No. gr-qc/9908061, 1999.

12. (136) S. Manoff, Deviation equations of Synge and Schild over spaces with affine connections and metrics, Intern. J. Mod. Phys. A, vol. 16, No. 6, pp. 1109-1122, 2001; http://arXiv.org e-Print archive, E-print No. gr-qc/0012015, 2000.

13. (137) S. Manoff, Mechanics of continuous media in (L̄_{n}, g) spaces. 1. Introduction and mathematical tools. http://arXiv.org e-Print archive, E-print No. gr-qc/0203016, March 2002.

14. (138) S. Manoff, B. Dimitrov, Flows and particles with shear - free and expansion - free velocities in (L̄_{n}, g) and weyl’s spaces, Classical and quantum gravity, vol. 19, pp. 4377-4398,2002 http://arXiv.org e-Print archive, E-print No. gr-qc/0207060, July 2002

15. (139) R. Aldrovandi, P. B. Barros, J. G. Pereira, The equivalence principle revisited, Foundations of physics, vol. 33, No. 4, pp. 545–575, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0212034, 2002.

16. (140) S. Manoff, B. Dimitrov, On the existence of a gyroscope in spaces with affine connections and metrics, General Relativity and Gravitation, vol. 35, pp. 25–33, 2003, http://arXiv.org e-Print archive, E-print No. gr-qc/0012011, 2000.

17. (141) S. Manoff, Deviation operator and deviation equations over spaces with affine connections and metrics, Journal of Geometry and Physics, Vol. 39, No. 4, October 2001, pp. 337-350.

18. (142) R. Aldrovandi, P. B. Barros, J. G. Pereira, Gravitation as anholonomy, General Relativity and Gravitation, vol. 35, No. 6, pp. 991–1005, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0301077, 2003.

19. (143) S. Manoff, Doppler effect and Hubble effect in different models of space-time in the case of auto-parallel motion of the observer, http://arXiv.org e-Print archive, E-print No. gr-qc/0401012, January 2004.

20. (144) Sawa Manoff, Geometry and mechanics in different models of space-time: Geometry and kinematics, Nova science Publishers, Inc., Yew York, 2002

21. (145) Sawa Manoff, Geometry and mechanics in different models of space-time: Dynamics and applications, Nova science Publishers, Inc., Yew York, 2002

22. (146) S. Manoff, Determination of the velocity of an emitter in spaces with affine connections and metrics, In “Applications and development of Lobachevski ideas in modern physics”, Proceedings of the International seminar dedicated to the 75th birthday of professor Nikolai A. Chernikov, Dubna 25–27 February 2004, Dubna, 2004. http://arXiv.org e-Print archive, E-print No. gr-qc/0403025, March 2004.

23. (147) S. Manoff, Standard emitters (clocks) and calibrated standard emitters (clocks) in spaces with affine connections and metrics, http://arXiv.org e-Print archive, E-print No. gr-qc/0505061, May 2005.

24. (148) Hunter Monroe, The equivalence principle along curves, local inflation, and nonsingular collapse (Second version: The Duality of Time Dilation and Velocity), http://arXiv.org e-Print archive, E-print No. gr-qc/0512019, December 2005.

25. (149) R. Aldrovandi, A. L. Barbosa, Connection space approach to ambiguities of gauge theories, International Journal of Mathematics and Mathematical Sciences, Vol. 2005, No. 15, pp. 2365–2371, 2005.

26. (150) Aldrovandi R. and J. G. Pereira, An Introduction to Teleparallel gravity, Instituto de Fisica Teorica, Universidade Estadual Paulista, Sao Paulo, Brazil, http://www.ift.unesp.br/gcg/tele.pdf, 2010.

27. (151) Giglio J. F. T. and W. A. Rodrigues Jr, Locally Inertial Reference Frames in Lorentzian and Riemann-Cartan Spacetimes, Analen der physik, Vol. 524, Issue 5, pp. 302–310, 2012, DOI: 10.1002/andp.201100302. http://arXiv.org e-Print archive, E-print No. arXiv:1111.2206v2 [math-ph], November 2011.

28. (152) Miguel Socolovsky, Einstein-Cartan Theory and Gauge Symmetry, http://arXiv.org e-Print archive, E-print No. arXiv:1212.3266 [gr-qc], December 2012.

29. (153) Aldrovandi R. and J. G. Pereira, An Introduction to teleparallel gravity, Fundamental Theories of Physics, Vol. 173, pp. 179–186, 2013, DOI 10.1007/978-94-007-5143-9_18.

30. (154) R. Aldrovandi, J. G. Pereira, An Introduction to teleparallel gravity, Instituto de Fisica Teorica, Universidade Estadual Paulista, Sao Paulo, Brazil, 2013.

31. (155) Tzanetatos Dimitrios, Determination of the ratio M/R via numerical solution the Tolman-Oppenheimer-Volkov equation for non-rotating spherically symmetric neutron stars and white dwarfs in the context of general relativity (in Greek), Bachelor’s Thesis, Supervisor associate professor Koutsoubas George, National Technical University of Athens, Faculty of applyed mathematics and natural sciences, Athens, 2014.

The work [20] is cited by:

1. (156) S. Manoff, A. Kolarov, B. Dimitrov, (\overline{L}_{n}, g)-spaces. General relativity over \overline{V}_{4}- spaces, Communication JINR, E5-98-184, Dubna, 1998.

2. (157) Vesselin Petkov, Does gravitational redshift support the curved-spacetime interpretation of general relativity?, http://arXiv.org e-Print archive, E-print No. gr-qc/9810030, 1998.

3. (158) S. Manoff, Relative velocity and relative acceleration induced by the torsion in (pseudo) Riemannian spaces with torsion and in spaces with an affine connection and metrics, http://arXiv.org e-Print archive, E-print No. gr-qc/0005027, 2000.

4. (159) S. Manoff, Spaces with contravariant and covariant affine connections and metrics, Part. Nucl. vol. 30, 517-549, 1999; Fiz. Elem. Chast. Atom. Yadra vol. 30, No. 3 1211-1269, 1999. http://arXiv.org e-Print archive, E-print No. gr-qc/0006024, 2000.

5. (160) S. Manoff, Auto-parallel equation as Euler-Lagrange’s equation in spaces with affine connections and metrics, Gen. Rel. Grav., vol. 32, No. 8, pp. 1559–1582, 2000, http://arXiv.org e-Print archive, E-print No. gr-qc/0010048, 2000.

6. (161) S. Manoff, Frames of reference in spaces with affine connections and metrics, Classical and quantum gravity, vol. 18, No. 6, 2001, pp. 1111–1126. http://arXiv.org e-Print archive, E-print No. gr-qc/9908061, 1999.

7. (162) S. Manoff, B. Dimitrov, Weyl’s spaces with shear-free and expansion-free conformal Killing vectors and the motion of a free spinless test particle, http://arXiv.org e-Print archive, E-print No. gr-qc/0011045, 2000.

8. (163) S. Manoff, Deviation operator and deviation equations over spaces with affine connections and metrics, Journal of Geometry and Physics, Vol. 39, No. 4, October 2001, pp. 337-350.

9. (164) S. Manoff, Deviation equations of Synge and Schild over spaces with affine connections and metrics, Intern. J. Mod. Phys. A, vol. 16, No. 6, pp. 1109-1122, 2001; http://arXiv.org e-Print archive, E-print No. gr-qc/0012015, 2000.

10. (165) Jakub Czajko, On conjugate complex time III: Superstrings and complex Lorentz transformation, Chaos, Solitons & Fractals , Vol 12, No 5 , 2001, pp. 951-967

11. (166) S. Manoff, Mechanics of continuous media in (L̄_{n}, g) spaces. 1. Introduction and mathematical tools. http://arXiv.org e-Print archive, E-print No. gr-qc/0203016, March 2002.

12. (167) R. Aldrovandi, P. B. Barros, J. G. Pereira, The equivalence principle revisited, Foundations of physics, vol. 33, No. 4, pp. 545–575, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0212034, 2002.

13. (168) S. Manoff, B. Dimitrov, On the existence of a gyroscope in spaces with affine connections and metrics, General Relativity and Gravitation, vol. 35, pp. 25–33, 2003, http://arXiv.org e-Print archive, E-print No. gr-qc/0012011, 2000.

14. (169) Sawa Manoff, Geometry and mechanics in different models of space-time: Geometry and kinematics, Nova science Publishers, Inc., Yew York, 2002

15. (170) Sawa Manoff, Geometry and mechanics in different models of space-time: Dynamics and applications, Nova science Publishers, Inc., Yew York, 2002

16. (171) S. Manoff, B. Dimitrov, Flows and particles with shear - free and expansion - free velocities in (L̄_{n}, g) and weyl’s spaces, Classical and quantum gravity, vol. 19, pp. 4377-4398,2002 http://arXiv.org e-Print archive, E-print No. gr-qc/0207060, July 2002

17. (172) R. Aldrovandi, P. B. Barros, J. G. Pereira, Gravitation as anholonomy, General Relativity and Gravitation, vol. 35, No. 6, pp. 991–1005, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0301077, 2003.

18. (173) C. Y. Lo, On Criticisms of Einstein’s Equivalence Principle, Physics Essays, Volume 16, No.1,March 2003.

19. (174) S. Manoff, Standard emitters (clocks) and calibrated standard emitters (clocks) in spaces with affine connections and metrics, http://arXiv.org e-Print archive, E-print No. gr-qc/0505061, May 2005.

20. (175) David Hestenes, Spacetime Geometry with Geometric Calculus, Document http://modelingnts.la.asu.edu/pdf/SpacetimeGeometry.w.GC.proc.pdf, Preceedings of the Seventh International Conference on Clifford Algebra, 2005.

21. (176) Giglio J. F. T. and W. A. Rodrigues Jr, Locally Inertial Reference Frames in Lorentzian and Riemann-Cartan Spacetimes, Analen der physik, Vol. 524, Issue 5, pp. 302–310, 2012, DOI: 10.1002/andp.201100302. http://arXiv.org e-Print archive, E-print No. arXiv:1111.2206v2 [math-ph], November 22. (177) R. Aldrovandi, J. G. Pereira, An Introduction to teleparallel gravity, Instituto

de Fisica Teorica, Universidade Estadual Paulista, Sao Paulo, Brazil, 2013.

23. (178) Enzo Zanchini, Reply to Comment on �Correct interpretation of two experiments on the transverse Doppler shift�, Physica Scripta, Volume 89, Issue 6, article id. 067005, DOI: 10.1088/0031-8949/89/6/067005, 2014.

24. (179) James Read and Nicholas J. Teh, Newtonian Equivalence Principles, Erkenntnis, Springer, 2022, DOI: 10.1007/s10670-021-00513-7, https://doi.org/10.1007/s10670-021-00513-7

The work [21] is cited by:

1. (180) Roman Plyatsko, Gravitational ultrarelativistic spin orbit interaction and the weak equivalence principle, Phys.Rev.D, vol. 5808, p. 4031, 1998.

2. (181) S. Manoff, A. Kolarov, B. Dimitrov, (\overline{L}_{n}, g)-spaces. General relativity over \overline{V}_{4}- spaces, Communication JINR, E5-98-184, Dubna, 1998.

3. (182) S. Manoff, Relative velocity and relative acceleration induced by the torsion in (pseudo) Riemannian spaces with torsion and in spaces with an affine connection and metrics, http://arXiv.org e-Print archive, E-print No. gr-qc/0005027, 2000.

4. (183) S. Manoff, Spaces with contravariant and covariant affine connections and metrics, Part. Nucl. vol. 30, 517-549, 1999; Fiz. Elem. Chast. Atom. Yadra vol. 30, No. 3 1211-1269, 1999. http://arXiv.org e-Print archive, E-print No. gr-qc/0006024, 2000.

5. (184) S. Manoff, Auto-parallel equation as Euler-Lagrange’s equation in spaces with affine connections and metrics, Gen. Rel. Grav., vol. 32, No. 8, pp. 1559–1582, 2000, http://arXiv.org e-Print archive, E-print No. gr-qc/0010048, 2000.

6. (185) S. Manoff, Frames of reference in spaces with affine connections and metrics, Classical and quantum gravity, vol. 18, No. 6, 2001, pp. 1111–1126. http://arXiv.org e-Print archive, E-print No. gr-qc/9908061, 1999.

7. (186) S. Manoff, B. Dimitrov, Weyl’s spaces with shear-free and expansion-free conformal Killing vectors and the motion of a free spinless test particle, http://arXiv.org e-Print archive, E-print No. gr-qc/0011045, 2000.

8. (187) S. Manoff, Deviation equations of Synge and Schild over spaces with affine connections and metrics, Intern. J. Mod. Phys. A, vol. 16, No. 6, pp. 1109-1122, 2001; http://arXiv.org e-Print archive, E-print No. gr-qc/0012015, 2000.

9. (188) S. Manoff, Mechanics of continuous media in (L̄_{n}, g) spaces. 1. Introduction and mathematical tools. http://arXiv.org e-Print archive, E-print No. gr-qc/0203016, March 2002.

10. (189) R. Aldrovandi, P. B. Barros, J. G. Pereira, The equivalence principle revisited, Foundations of physics, vol. 33, No. 4, pp. 545–575, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0212034, 2002.

11. (190) S. Manoff, B. Dimitrov, On the existence of a gyroscope in spaces with affine connections and metrics, General Relativity and Gravitation, vol. 35, pp. 25–33, 2003, http://arXiv.org e-Print archive, E-print No. gr-qc/0012011, 2000.

12. (191) Sawa Manoff, Geometry and mechanics in different models of space-time: Geometry and kinematics, Nova science Publishers, Inc., Yew York, 2002

13. (192) S. Manoff, Deviation operator and deviation equations over spaces with affine connections and metrics, Journal of Geometry and Physics, Vol. 39, No. 4, October 2001, pp. 337-350.

14. (193) Sawa Manoff, Geometry and mechanics in different models of space-time: Dynamics and applications, Nova science Publishers, Inc., Yew York, 2002

15. (194) S. Manoff, B. Dimitrov, Flows and particles with shear - free and expansion - free velocities in (L̄_{n}, g) and weyl’s spaces, Classical and quantum gravity, vol. 19, pp. 4377-4398,2002 http://arXiv.org e-Print archive, E-print No. gr-qc/0207060, July 2002

16. (195) R. Aldrovandi, P. B. Barros, J. G. Pereira, Gravitation as anholonomy, General Relativity and Gravitation, vol. 35, No. 6, pp. 991–1005, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0301077, 2003.

17. (196) S. Manoff, Doppler effect and Hubble effect in different models of space-time in the case of auto-parallel motion of the observer, http://arXiv.org e-Print archive, E-print No. gr-qc/0401012, January 2004.

18. (197) S. Manoff, Determination of the velocity of an emitter in spaces with affine connections and metrics, In “Applications and development of Lobachevski ideas in modern physics”, Proceedings of the International seminar dedicated to the 75th birthday of professor Nikolai A. Chernikov, Dubna 25–27 February 2004, Dubna, 2004. http://arXiv.org e-Print archive, E-print No. gr-qc/0403025, March 2004.

19. (198) R. Aldrovandi, A. L. Barbosa, Connection space approach to ambiguities of gauge theories, International Journal of Mathematics and Mathematical Sciences, Vol. 2005, No. 15, pp. 2365–2371, 2005.

20. (199) Aldrovandi R. and J. G. Pereira, An Introduction to Teleparallel gravity, Instituto de Fisica Teorica, Universidade Estadual Paulista, Sao Paulo, Brazil, http://www.ift.unesp.br/gcg/tele.pdf, 2010.

21. (200) Giglio J. F. T. and W. A. Rodrigues Jr, Locally Inertial Reference Frames in Lorentzian and Riemann-Cartan Spacetimes, Analen der physik, Vol. 524, Issue 5, pp. 302–310, 2012, DOI: 10.1002/andp.201100302. http://arXiv.org e-Print archive, E-print No. arXiv:1111.2206v2 [math-ph], November 2011.

22. (201) Andrew J.S. Hamilton and Pedro P. Avelino, The Physics of the relativistic counter-streaming instability that drives mass inflation inside black holes, Phys.Rept. 495, pp. 1-32, 2010, http://arXiv.org e-Print archive, E-print No. arXiv:0811.1926 [gr-qc], 2010.

23. (202) Aldrovandi R. and J. G. Pereira, An Introduction to teleparallel gravity, Fundamental Theories of Physics, Vol. 173, pp. 179–186, 2013, DOI 10.1007/978-94-007-5143-9_18.

24. (203) R. Aldrovandi, J. G. Pereira, An Introduction to teleparallel gravity, Instituto de Fisica Teorica, Universidade Estadual Paulista, Sao Paulo, Brazil, 2013.

The work [22] is cited by:

1. (204) A. Echeverria-Enriquez, M. C. Munoz-Lecanda, N. Roman-Roy, C. Victoria-Monge, Mathematical Foundations of Geometric Quantization, Extracta Math., vol. 13, pp. 135-238, 1998. http://arXiv.org e-Print archive, E-print No. math-ph/9904008, 1999.

2. (205) G. Sardanashvily, On quantum evolution as a parallel transport, http://arXiv.org e-Print archive, E-print No. quant-ph/0004050, 2000.

3. (206) L. Mangiarotti, G. Sardanashvily, Connections in classical and quantum field theory, World Scientific, Singapore-New Jersey-London-Hong Kong, 2000. (Reference No. 160.)

4. (207) Pravabati Chingangbam, Pankaj Sharan, Pseudoforces in quantum mechanics, Physical Review A, vol. 64, No. 042107, 2001. http://arXiv.org e-Print archive, E-print No. quant-ph/0105074, 2001.

The work [23] is cited by:

1. (208) Reza Mansouri, Kourosh Nozari, Signature Change, Inflation, and the Cosmological Constant, Preprint IPM/P-98/15, 1998. http://arXiv.org e-Print archive, E-print No. gr-qc/9806109, 1998.

2. (209) Reza Mansouri, Kourosh Nozari, Application of Colombeau’s Generalized Functions to Cosmological Models with Signature Change, Talk given at International Seminar on Mathematical Cosmology (ISMC 98), Potsdam, Germany, 30 Mar - 4 Apr 1998. Preprint IPM-98-283. http://arXiv.org e-Print archive, E-print No. gr-qc/9806007, 1998.

3. (210) Reza Mansouri, Kourosh Nozari, A New Distributional Approach to Signature Change, General Relativity and Gravitation, Vol. 32, No. 2, pp. 253–269, 2000.

4. (211) K Nozari, New Junction Conditions for Signature Change: Null Boundary Proposal, International Journal of Theoretical Physics, vol. 44, Number 2, pp. 235-244, February, 2005

5. (212) Craig Callender, Time is the Simplest (and Strongest) Thing, Centre for Time, Department of Philosophy, University of Sydney, 2006, http://ses.library.usyd.edu.au/handle/2123/1071, http://ses.library.usyd.edu.au/bitstream/2123/1071/2/Callender1.ppt

The work [?] is cited by:

1. (213) G. Sardanashvily, On quantum evolution as a parallel transport, http://arXiv.org e-Print archive, E-print No. quant-ph/0004050, 2000.

2. (214) Tony Rothman, George F. R. Ellis, Jeff Murugan, Holonomy in the Schwarzschild-Droste Geometry, Class. Quant. Grav. vol 18, pp. 1217-1234, 2001; http://arXiv.org e-Print archive, E-print No. gr-qc/0008070, 2000.

3. (215) D. D. Ferrante Geometric calculus and the fibre bundle description of quantum mechanics, Preprint BROWN-HET-1307, March 2002; http://arXiv.org e-Print archive, E-print No. math-ph/0204024, 2002.

4. (216) G.Giachetta, L.Mangiarotti, G.Sardanashvily, Geometric quantization of completely integrable Hamiltonian systems in the action-angle variables, Physics letters A, vol. 301, pp. 53–57, 2002. http://arXiv.org e-Print archive, E-print No. quant-ph/0112083, 2001.

5. (217) E.Fiorani, G.Giachetta, G.Sardanashvily, Geometric quantization of time-dependent completely integrable Hamiltonian systems, J. Math. Phys. 43 (2002) 5013-5025. http://arXiv.org e-Print archive, E-print No. quant-ph/0202093, 2002.

6. (218) Daniel D. Ferrante, Generalized Bundle Quantum Mechanics, Preprint BROWN-HET-1374, 2003; http://arXiv.org e-Print archive, E-print No. math-ph/0308004, August 2003.

7. (219) G.Giachetta, L.Mangiarotti, G.Sardanashvily, Non-adiabatic holonomy operators in classical and quantum completely integrable systems, Journal of mathematical physics, vol. 45, No. 1, pp. 76–86, 2004. http://arXiv.org e-Print archive, E-print No. quant-ph/0212108, 2002.

8. (220) G. Sardanashvily, G. Giachetta, What is geometry in quantum theory, International Journal of Geometric Methods in Modern Physics, vol. I, No. 1 & 2, pp. 1–22, 2004. http://arXiv.org e-Print archive, E-print No. hep-th/0401080, 2004.

9. (221) David Viennot, Principal bundle structure of quantum adiabatic dynamics with a Berry phase which does not commute with the dynamical phase, Journal of Mathematical Physics, vol. 46, No. 072102, 2005

10. (222) David Viennot, Géométrie et Adiabaticité des Systémes Photodynamiques Quantiques, These de Doctorat SpćialitṔhysique, Université de Franche-Comé, Ecole doctorale Louis Pasteur, 2005.

11. (223) Katarzyna Grabowska, Janusz Grabowski and Pawel Urbański, The Schroedinger operator in Newtonian space-time, http://arXiv.org e-Print archive, E-print No. math-ph/0611044, 2006.

12. (224) Paul Benioff, Effects of a scalar scaling field on quantum mechanics, http://arXiv.org e-Print archive, E-print No. arXiv:1512.05669 [quant-ph], 2015.

13. (225) Paul Benioff, Effect of number scaling on entangled states in quantum mechanics, Proc. SPIE Int. Soc. Opt. Eng. 9873 (2016) 98730D; Proceedings Volume 9873, Quantum Information and Computation IX; 98730D (2016), https://doi.org/10.1117/12.2221958, Event: SPIE Commercial + Scientific Sensing and Imaging, 2016, Baltimore, Maryland, United States http://arXiv.org e-Print archive, E-print No. arXiv:1603.04752 [quant-ph], 2016.

14. (226) Thomas Bittner, Towards a Quantum Theory of Geographic Fields, 13th International Conference on Spatial Information Theory (COSIT 2017), Editors: Eliseo Clementini, Maureen Donnelly, May Yuan, Christian Kray, Paolo Fogliaroni, and Andrea Ballatore; Article No. 5; pp. 5:1�5:14, Leibniz International Proceedings in Informatics, Schloss Dagstuhl � Leibniz-Zentrum fьr Informatik, Dagstuhl Publishing, Germany, 2017.

The work [24] is cited by:

1. (227) A. Echeverria-Enriquez, M. C. Munoz-Lecanda, N. Roman-Roy, C. Victoria-Monge, Mathematical Foundations of Geometric Quantization, Extracta Math. vol. 13, pp. 135-238, 1998. http://arXiv.org e-Print archive, E-print No. math-ph/9904008, 1999.

2. (228) Vadim G. Zhotikov, Geometry of variational analysis and its applications to theoretical physics, Science and technology litrerature Publisher, Tomsk, 2002, In Russian.

3. (229) G.Giachetta, L.Mangiarotti, G.Sardanashvily, Geometric quantization of completely integrable Hamiltonian systems in the action-angle variables, Physics letters A, vol. 301, pp. 53–57, 2002. http://arXiv.org e-Print archive, E-print No. quant-ph/0112083, 2001.

4. (230) E.Fiorani, G.Giachetta, G.Sardanashvily, Geometric quantization of time-dependent completely integrable Hamiltonian systems, J. Math. Phys. 43 (2002) 5013-5025. http://arXiv.org e-Print archive, E-print No. quant-ph/0202093, 2002.

5. (231) G.Giachetta, L.Mangiarotti, G.Sardanashvily, Non-adiabatic holonomy operators in classical and quantum completely integrable systems, Journal of mathematical physics, vol. 45, No. 1, pp. 76–86, 2004. http://arXiv.org e-Print archive, E-print No. quant-ph/0212108, 2002.

6. (232) G. Sardanashvily, G. Giachetta, What is geometry in quantum theory, International Journal of Geometric Methods in Modern Physics, vol. I, No. 1 & 2, pp. 1–22, 2004. http://arXiv.org e-Print archive, E-print No. hep-th/0401080, 2004.

7. (233) G.Giachetta, L.Mangiarotti, G.Sardanashvily, Geometric and Algebraic Topological Methods in Quantum Mechanics, World Scientific, Singapore, 2005. http://arXiv.org e-Print archive, E-print No. math-ph/0410040, 2004.

8. (234) David Viennot, Principal bundle structure of quantum adiabatic dynamics with a Berry phase which does not commute with the dynamical phase, Journal of Mathematical Physics, vol. 46, No. 072102, 2005.

9. (235) David Viennot, Géométrie et Adiabaticité des Systémes Photodynamiques Quantiques, These de Doctorat SpćialitṔhysique, Université de Franche-Comé, Ecole doctorale Louis Pasteur, 2005.

10. (236) Katarzyna Grabowska, Janusz Grabowski and Pawel Urbański, The Schroedinger operator in Newtonian space-time, http://arXiv.org e-Print archive, E-print No. math-ph/0611044, 2006.

11. (237) Katarzyna Grabowska, Janusz Grabows and Pawel Urbanśki The Schroedinger operator as a generalized Laplacian, J. Phys. A: Math. Theor, vol. 41, 145204. 2008, DOI: 10.1088/1751-8113/41/14/145204

12. (238) D. J. Hurley and M. Vandyck, D-differentiation in Hilbert space and the structure of Quantum Mechanics, Foundations of Physics, vol. 30, pp.433-473 2009.

13. (239) D. J. Hurley and M. Vandyck, D-differentiation in Hilbert space and the structure of Quantum Mechanics, II. Accelerated observers and fictitious forces. Foundations of Physics vol. 39, No. 5, pp. 433-473, 2009, DOI: 10.1007/s10701-009-9297-6.

14. (240) D. J. Hurley and M. A. Vandyck, Differentiation in Hilbert Space and the Structure of Quantum Mechanics Part II: Accelerated Observers and Fictitious Forces, Foundations of Physics vol. 41, No. 4 , pp. 667–685, 2011.

15. (241) Liang Dong and Qian Niu, Geometrodynamics of electrons in a crystal under position and time dependent deformation, http://arXiv.org e-Print archive, E-print No. arXiv:1802.02887, 8 February 2018.

16. (242) Marcelo Epstein, Hilbert bundles as quantum-classical continua, Mathematics and Mechanics of Solids, Vol 25, Issue 6, pp. 1312�1317, June 17 2020, DOI: 10.1177/1081286519888964. http://arXiv.org e-Print archive, E-print No. arXiv:1908.09069 [quant-ph], 2019.

17. (243) Hoi Van Nguyen, Loпc Le Marrec and Guy Casale, On tangent geometry and generalised continuum with defects, Reference (Citation) No.44, Preprint in Mathematics and Mechanics of Solids, November 2021, DOI: 10.1177/10812865211059222 Foundations of Physics, vol. 30, pp.433-473, 2009.

18. (244) Varga Bonbien Aurelien and Manchon Aurelien Manchon, Theory of perturbatively nonlinear quantum transport II: Hilbert space truncation, gauge invariance, and second order transport in a spatially uniform, time-varying electric field, http://arXiv.org e-Print archive, E-print No. 2202.05378 [cond-mat.mes-hall], February 2022, https://doi.org/10.48550/arXiv.2202.05378; Citation No.41.

19. (245) Hoi Van Nguyen, Geometric models and applications to material media with defects, Doctoral Thesis, Universi’e Rennes 1, Advisor Loic Le Marrec, 2022, http://www.theses.fr/s209162, https://www.researchgate.net/publication/358893809_Geometric_models_and_applications_to_material_media_with_defects See pages 82 and 130.

The work [25] is cited by:

1. (246) M. Reuter, Symplectic Dirac-Kähler Fields, J. Math. Phys., vol. 40, pp. 5593-5640, 1999. http://arXiv.org e-Print archive, E-print No. hep-th/9910085, 1999.

2. (247) Vadim G. Zhotikov, Geometry of variational analysis and its applications to theoretical physics, Science and technology litrerature Publisher, Tomsk, 2000, In Russian.

3. (248) G. Sardanashvily, G. Giachetta, What is geometry in quantum theory, International Journal of Geometric Methods in Modern Physics, vol. I, No. 1 & 2, pp. 1–22, 2004. http://arXiv.org e-Print archive, E-print No. hep-th/0401080, 2004.

4. (249) G.Giachetta, L.Mangiarotti, G.Sardanashvily, Geometric and Algebraic Topological Methods in Quantum Mechanics, World Scientific, Singapore, 2005. http://arXiv.org e-Print archive, E-print No. math-ph/0410040, 2004.

5. (250) David Viennot, Principal bundle structure of quantum adiabatic dynamics with a Berry phase which does not commute with the dynamical phase, Journal of Mathematical Physics, vol. 46, No. 072102, 2005

6. (251) Katarzyna Grabowska, Janusz Grabowski and Pawel Urbański, The Schroedinger operator in Newtonian space-time, http://arXiv.org e-Print archive, E-print No. math-ph/0611044, 2006.

7. (252) Katarzyna Grabowska, Janusz Grabows and Pawel Urba The Schroedinger operator as a generalized Laplacian, J. Phys. A: Math. Theor, vol. 41, 145204. 2008, DOI: 10.1088/1751-8113/41/14/145204

8. (253) D. J. Hurley and M. Vandyck, D-differentiation in Hilbert space and the structure of Quantum Mechanics, Foundations of Physics, vol. 30, pp.433-473, 2009.

9. (254) Varga Bonbien and Aurelien Manchon, Theory of perturbatively nonlinear quantum transport II: Hilbert space truncation, gauge invariance, and second order transport in a spatially uniform, time-varying electric field, http://arXiv.org e-Print archive, E-print No. 2202.05378 [cond-mat.mes-hall], February 2022, https://doi.org/10.48550/arXiv.2202.05378 Citation No.43.

The work [26] is cited by:

1. (255) Vadim G. Zhotikov, Geometry of variational analysis and its applications to theoretical physics, Science and technology litrerature Publisher, Tomsk, 2000, In Russian.

2. (256) G. Sardanashvily, G. Giachetta, What is geometry in quantum theory, International Journal of Geometric Methods in Modern Physics, vol. I, No. 1 & 2, pp. 1–22, 2004. http://arXiv.org e-Print archive, E-print No. hep-th/0401080, 2004.

3. (257) G.Giachetta, L.Mangiarotti, G.Sardanashvily, Geometric and Algebraic Topological Methods in Quantum Mechanics, World Scientific, Singapore, 2005. http://arXiv.org e-Print archive, E-print No. math-ph/0410040, 2004.

4. (258) David Viennot, Principal bundle structure of quantum adiabatic dynamics with a Berry phase which does not commute with the dynamical phase, Journal of Mathematical Physics, vol. 46, No. 072102, 2005

5. (259) Katarzyna Grabowska, Janusz Grabowski and Pawel Urbański, The Schroedinger operator in Newtonian space-time, http://arXiv.org e-Print archive, E-print No. math-ph/0611044, 2006.

6. (260) Katarzyna Grabowska, Janusz Grabows and Pawel Urbanśki The Schroedinger operator as a generalized Laplacian, J. Phys. A: Math. Theor, vol. 41, 145204. 2008, DOI: 10.1088/1751-8113/41/14/145204

7. (261) D. J. Hurley and M. Vandyck, D-differentiation in Hilbert space and the structure of Quantum Mechanics, Foundations of Physics, vol. 30, pp.433-473, 2009.

8. (262) Marcelo Epstein, Hilbert bodies as quantum-classical continua, http://arXiv.org e-Print archive, E-print No. 1908.09069 [quant-ph], 2019.

9. (263) Gao Zhi Fu1, Chen Ci Xing and Wang Na, The scattering of Dirac spinors in rotating spheroids, The European Physical Journal C, 80:582, 2020, DOI: 10.1140/epjc/s10052-020-8140-2. Foundations of Physics, vol. 30, pp.433-473, 2009.

10. (264) Varga Bonbien and Aurelien Manchon, Theory of perturbatively nonlinear quantum transport II: Hilbert space truncation, gauge invariance, and second order transport in a spatially uniform, time-varying electric field, http://arXiv.org e-Print archive, E-print No. 2202.05378 [cond-mat.mes-hall], February 2022, https://doi.org/10.48550/arXiv.2202.05378, Citation No.44.

The work [27] is cited by:

1. (265) D. J. Hurley and M. Vandyck, D-differentiation in Hilbert space and the structure of Quantum Mechanics, Foundations of Physics, vol. 30, pp.433-473, 2009.

2. (266) Pankaj Sharan, A Bundle Picture of Quantum Mechanics, In: Some Unusual Topics in Quantum Mechanics. Springer Briefs in Physics. Springer, 2020, DOI https://doi.org/10.1007/978-3-030-60418-9.

The work [28] is cited by:

1. (267) D. J. Hurley and M. Vandyck, D-differentiation in Hilbert space and the structure of Quantum Mechanics, Foundations of Physics, vol. 30, pp.433-473, 2009.

2. (268) D. J. Hurley and M. Vandyck, D-differentiation in Hilbert space and the structure of Quantum Mechanics, II. Accelerated observers and fictitious forces. Foundations of Physics vol. 39, No. 5, pp. 433-473, 2009, DOI: 10.1007/s10701-009-9297-6.

3. (269) D. J. Hurley and M. A. Vandyck, Differentiation in Hilbert Space and the Structure of Quantum Mechanics Part II: Accelerated Observers and Fictitious Forces, Foundations of Physics vol. 41, No. 4 , pp. 667–685, 2011.

4. (270) Pankaj Sharan, A Bundle Picture of Quantum Mechanics, In: Some Unusual Topics in Quantum Mechanics. Series Lecture Notes in Physics. Springer, 2023, DOI https://doi.org/10.1007/978-3-031-35962-0.

Softcover ISBN 978-3-031-35961-3, Published: 22 September 2023. eBook ISBN 978-3-031-35962-0, Published: 21 September 2023. Edition Number 2.

The work [29] is cited by:

1. (271) Reza Mansouri, Kourosh Nozari, Signature Change, Inflation, and the Cosmological Constant, Preprint IPM/P-98/15, 1998. http://arXiv.org e-Print archive, E-print No. gr-qc/9806109, 1998.

2. (272) Reza Mansouri, Kourosh Nozari, Application of Colombeau’s Generalized Functions to Cosmological Models with Signature Change, Talk given at International Seminar on Mathematical Cosmology (ISMC 98), Potsdam, Germany, 30 Mar - 4 Apr 1998. Preprint IPM-98-283. http://arXiv.org e-Print archive, E-print No. gr-qc/9806007, 1998.

3. (273) K Nozari, New Junction Conditions for Signature Change: Null Boundary Proposal, International Journal of Theoretical Physics, vol. 44, Number 2, pp. 235-244, February, 2005

4. (274) Duplij Stepsn Semigroup Methods In Supersymmetric Theories Of Elementary Particles, Habilitation Thesis in Russian, 483 pp (Kharkov State University, Ukraine). http://arXiv.org e-Print archive, E-print No. math-ph/9910045, 1999. (Reference No. 197.)

5. (275) Duplij S. A., Semisupermenifolds and semigroups, “Krok” Publisher, Kharkov, 2000. (Reference No. 553.)

6. (276) Reza Mansouri, Kourosh Nozari, A New Distributional Approach to Signature Change, General Relativity and Gravitation, Vol. 32, No. 2, pp. 253–269, 2000.

7. (277) A. Das, A. DeBenedictis, N. Tariq, General solutions of Einstein’s spherically symmetric gravitational equations with junction conditions, J. Math. Phys. vol. 44, 2003, pp. 5637-5655. http://arXiv.org e-Print archive, E-print No. gr-qc/0307009, 2003.

8. (278) R.A. El-Nabulsi Nonstandard fractional exponential Lagrangians, fractional geodesic equation, complex general relativity, and discrete gravity, Canadian Journal of Physics, Volume 91, Issue 8, August 2013, Pages 618-622.

9. (279) Rami Ahmad El-Nabulsi, From Classical to Discrete Gravity through Exponential Non-Standard Lagrangians in General Relativity, Mathematics Volume 3, Issue 3, pp. 727-745, 2015; doi:10.3390/math3030727, http://www.mdpi.com/2227-7390/3/3/727/htm

The work [30] is cited by:

1. (280) D. D. Ferrante Geometric calculus and the fibre bundle description of quantum mechanics, Preprint BROWN-HET-1307, March 2002; http://arXiv.org e-Print archive, E-print No. math-ph/0204024, 2002.

2. (281) Daniel D. Ferrante, Generalized Bundle Quantum Mechanics, Preprint BROWN-HET-1374, 2003; http://arXiv.org e-Print archive, E-print No. math-ph/0308004, August 2003.

The work [31] is cited by:

1. (282) E. B. Manoukian, S. Sukkhasena and S. Siranan, Variational derivatives of transformation functions in quantum field theory, Physica Scripta, vol. 75, p. 751, 2007, DOI: 10.1088/0031-8949/75/6/001

2. (283) Sukkhasena Seckson, Closed�time path in quantum physics and quantum field theory, A Thesis for the Degree of Doctor of Philosophy in Physics Suranaree University of Technology, Academic Year 2007.

3. (284) E. B. Manoukian1 and K. Limboonsong, Quadratic Actions in Dependent Fields and the Action Principle, International Journal of Theoretical Physics, Volume 47, Number 5, pp. 1424-1431, May, 2008, DOI: 10.1007/s10773-007-9584-y

4. (285) E. B. Manoukian and K. Limboonsong , Constraints, dependent fields and the quantum dynamical principle, Physica Scripta, vol.77, 065010, 2008 DOI: 10.1088/0031-8949/77/06/065010

5. (286) Tukkamon Vijaktanawudhi, Quantum dynamical principle of constrained dynamics in quantum physics and quantum field theory, A Thesis Submitted in Partial Fulfillment of the Requirements for the Degree of Doctor of Philosophy, Physics Suranaree University of Technology Academic Year, 2008,

The work [32] is cited by:

1. (287) R. Aldrovandi, P. B. Barros, J. G. Pereira, The equivalence principle revisited, Foundations of physics, vol. 33, No. 4, pp. 545–575, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0212034, 2002.

2. (288) R. Aldrovandi, P. B. Barros, J. G. Pereira, Gravitation as anholonomy, General Relativity and Gravitation, vol. 35, No. 6, pp. 991–1005, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0301077, 2003.

3. (289) R. Aldrovandi, A. L. Barbosa, Connection space approach to ambiguities of gauge theories, International Journal of Mathematics and Mathematical Sciences, Vol. 2005, No. 15, pp. 2365–2371, 2005.

4. (290) James M. Nester, Normal frames for general connections, Ann. Phys. (Berlin), vol. 19, No. 1�2, pp. 45�52, 2010, DOI: 10.1002/andp.200910373

5. (291) Garecki Janusz, Is torsion needed in a theory of gravity? A reappraisal, http://arXiv.org e-Print archive, E-print No. arXiv:1110.4251v1 [gr-qc], October 2011.

6. (292) Ralitza K. Kovacheva, Montel’s Type Results and Zero Distribution of Sequences of rational Functions, Bulletin de la Societe des sciences et des letters de Lodz, Se�rie: Recherches sur les de�formations, In memory of Professor Roman Stanislaw Ingarden, Volume LXI, No. 3, pp. 7�-13, Lodz, 2011.

The work [33] is cited by:

1. (293) Abdelmoumene Belabbas, Les Interactions Fondamentales et la Structure de l’Espacetemps http://arXiv.org e-Print archive, E-print No. arXiv:1401.0800 [gr-qc], 2014.

2. (294) Alessio Rocci, La storia della Gravitа Quantistica: Dalla nascita della Relativitа Generale al secondo dopoguerra (1915-1945) (In Italian, History of Quantum Gravity: from the birth of General Relativity to the end of WWII (1915-1945)), Universitа degli Studi di Padova, Dipartimento di Fisica e Astronomia ”Galileo Galilei”, Ph.D. Theses, 297 pp, July 30, 2016.

The work [34] is cited by:

1. (295) Exercises of The Mathematica GuideBook for Symbolics, Document http://www.mathematicaguidebooks.org/scripts/download file.cgi?software download=Sample_Exercises_Symbolics.nb.pdf from the site www.mathematicaguidebooks.org.

2. (296) Notebooks accompanying The Mathematica GuideBook for Programming, Springer,

The work [35] is cited by:

1. (297) Kiyoumars A. Sohrabi, Microscopic Study of Vorticities in Relativistic Chiral Fermions, http://arXiv.org e-Print archive, E-print No. 1409.4528 [hep-ph], 2014.

The work [36] is cited by:

1. (298) R. Aldrovandi, P. B. Barros, J. G. Pereira, Gravitation as anholonomy, General Relativity and Gravitation, vol. 35, No. 6, pp. 991–1005, 2003; http://arXiv.org e-Print archive, E-print No. gr-qc/0301077, 2003.

The work [37] is cited by:

1. (299) Tobias Fritz, Lokalkompakte Gruppoide, University of Munster, e-Print URL: http://www.math.uni-muenster.de/u/rameyer/Morita_Gruppoide/Beispiele_Gruppoide.pdf, 15 April 2005

The work [38] is cited by:

1. (300) Daniel Arean and Alfonso V. Ramallo, Open string modes at brane intersections, JHEP 0604 (2006) 037, http://arXiv.org e-Print archive, E-print No. hep-th/0602174, February 2006.

2. (301) Daniel Arean Fraga, Holographic flavor in the gauge/gravity duality, http://arXiv.org e-Print archive, E-print No. arXiv:0906.5183 [hep-th], June 2009.

3. (302) Phillip E. Parker Lectures on Connection Geometry, 2013, Mathematics Department, Wichita State University, Wichita KS 67260-0033, USA, URL: http://www.math.wichita.edu/p̃parker/research/lcg.pdf

The work [39] is cited by:

1. (303) Carlo Cercignani, Slow Rarefied Flows: Theory and Application to Micro-Electro-Mechanical Systems, Progress in Mathematical Physics vol. 41, Birkhauser, 2006.

2. (304) Biological Physics: Poincarй Seminar 2009 (Progress in Mathematical Physics No.60), Birkhauser; 2011 edition (December 17, 2012), ISBN-10: 303480315X, ISBN-13: 978-3034803151 (See page 170).

3. (305) James M. Nester, Normal frames for general connections, Ann. Phys. (Berlin), vol. 19, No. 1�2, pp. 45�52, 2010, DOI: 10.1002/andp.200910373

4. (306) Jotz M., T. S. Ratiu and M. Zambon, Invariant frames for vector bundles and applications, Geometriae Dedicata, Springer, 2011, DOI: 10.1007/s10711-011-9618-8

5. (307) Jotz Madeleine, The leaf space of a multiplicative foliation, http://arXiv.org e-Print archive, E-print No. arXiv:1010.3127v3, v1 October 2010 and v3 October 2011

6. (308) Mintas Fatmagьl, The length function of a twistor spinor with zero, Annals of Global Analysis and Geometry, Vol. 41, Number 4, 493-514, Springer, 2011

7. (309) Knox Eleanor, Newton-Cartan theory and teleparallel gravity: the force of a formulation King�s College London, Pre-print, 20 September 2011.

8. (310) Christakos G., Space�Time and Uncertainty, Integrative problem-solving in a time of decadence, 2011, pp 205-242.

9. (311) Thomas M. and Muller Ewald, The Bondi-Sachs metric at the vertex of a null cone: axially symmetric vacuum solutions, Classical and Quantum Gravity, Volume 30, Number 5, p. 055019, 2013, DOI: 10.1088/0264-9381/30/5/055019; http://arXiv.org e-Print archive, E-print No. arXiv:1211.4980 [gr-qc], Nov. 2012 (v1) and Jan. 2013 (v2).

10. (312) Knox E., Effective spacetime geometry, Studies in history and philosophy of Science Part B: Studies in history and philosophy of modern physics, Vol. 44, Issue 3, August 2013, pp. 346�356, doi:10.1016/j.shpsb.2013.04.002.

11. (313) Milutin Blagojević and Friedrich W. Hehl (Editors) Gauge theories of gravitation. A reader with commentaries, Imperial College Press, London, April 2013 (See Chapter 5 General Structure of Poincar�e Gauge. Theory (Including Quadratic Lagrangians)).

12. (314) Bogdan Georgiev Dimitrov, Methods of algebraic geometry and Fermi-Walker transport in extended gravity theory (in Russian), Ph.D. Thesis,

Russian University Druzhba Narodov, Moskow, 2013.

13. (315) Homero G. Diaz-Marin, General Boundary Formulation for n-Dimensional Classical Abelian Theory with Corners, Symmetry, Integrability and Geometry: Methods and Applications (SIGMA), volume 11, 2015, DOI: 10.3842/SIGMA.2015.048, http://arXiv.org e-Print archive, E-print No. arXiv:1407.4741 [math-ph], v. 4, 2014.

14. (316) Homero G. Dнaz-Marнn, General Boundary Formulation for n-Dimensional Classical Abelian Theory with Corners, http://arXiv.org e-Print archive, E-print No. arXiv:1100592 [math-ph], 28 October 2014.

15. (317) Richard Chapling, A Hypergeometric Integral with Applications to the Fundamental Solution of Laplace�s Equation on Hyperspheres, http://arXiv.org e-Print archive, E-print No. arXiv:1508.06689 [math-ph], August 2015.

16. (318) Luca Rizzi and Pavel Silveira, Sub-Riemannian Ricci curvatures and universal diameter bounds for 3-Sasakian manifolds, http://arXiv.org e-Print archive, E-print No. arXiv:1509.05410 [math.DG], v. 2, October 2015.

17. (319) Crasmareanu, M. and Frigioiu, C., Unitary vector fields are Fermi-Walker transported along Rytov-Legendre curves, International Journal of Geometric Methods in Modern Physics, Volume 12, Issue 10, 1 November 2015, Article number 1550111, DOI: 10.1142/S021988781550111X.

18. (320) Homero G. Dнaz-Marin, General Boundary Formulation for n-Dimensional Classical Abelian Theory with Corners, Symmetry, Integrability and Geometry: Methods and Applications, SIGMA Volume 11, page 048, 35 pages, 2015.

19. (321) Homero Geovani Diaz Marin, Operador de Dirichlet-Neumann para campos abelianos de Yang-Mills en superficies, In Proceedings of the Second International Conference on Mathematics and its Applications su composiciуn, diseсo y cuidado, estuvo a cargo de Fernando Macнas Romero, Ediciуn electrуnica, 21.9 MB, disponible en, http://www.fcfm.buap.mx/cima/assets/docs/2015/memoriasCIMA2015.pdf

20. (322) R. Chapling, A Hypergeometric Integral with Applications to the Fundamental Solution of Laplace�s Equation on Hyperspheres, Symmetry, Integrability and Geometry: Methods and Applications (SIGMA), Volume 12, p. 079, 20 pages, 2016; http://dx.doi.org/10.3842/SIGMA.2016.079 21. (323) Davide Barilari and Stefan Ivanov, A Bonnet-Myers type theorem for quaternionic

contact structures, Calculus of Variations, vol. 58, No.1, Feb. 2019, http://arXiv.org e-Print archive, E-print No. arXiv:1703.04340 [math], 2017, 2018.

22. (324) Ivan Potashov and Alexander Tsiruev, Computational Algorithm for Covariant Series Expansions in General Relativity, The European Physical Journal Conferences, vol. 173(1):03021, January 2018, DOI: 10.1051/epjconf/201817303021.

23. (325) Ivan Potashov and Alexander Tsirulev, Computational Algorithm for Covariant Series Expansions in General Relativity, Faculty of Mathematics, Tver State University, Tver, Russia, EPJ Web Conf, Volume 173, 2018, Mathematical Modeling and Computational Physics 2017 (MMCP 2017), Article Number 03021, Number of pages 4, Section Numerical Modeling and Methods, DOI https://doi.org/10.1051/epjconf/201817303021, Published online 14 February 2018.

24. (326) Davide Barilari and Stefan Ivanov, A Bonnet-Myers type theorem for quaternionic contact structures, Calculus of Variations 58(1), DOI: 10.1007/s00526-018-1467-y, 2019.

25. (327) Naveed S. Bari and Eugenie Hunsicker, Isospectrality for Orbifold Lens Spaces, Canadian Journal of Mathematics, 1-45, DOI: 10.4153/S0008414X19000178, Cambridge University Press, 28 March 2019. http://arXiv.org e-Print archive, E-print No. arXiv:1705.01412 [math], 2017.

26. (328) Luca Rizzi and Pavel Silveira, Sub-Riemannian Ricci curvatures and universal diameter bounds for 3-Sasakian manifolds, Journal of the Institute of Mathematics of Jussieu, Volume 18, Issue 4, July 2019 , pp. 783–827, DOI: 10.1017/S1474748017000226, 2019; J. Inst. Math. Jussieu, pp. 1-45, 2017.

27. (329) Oliver Gabel, Bose-Einsteincondensation in curved space-time. From Concepts of General Relativity to Tidal Corrections for Quantum Gases in Local Frames, Dissertation, Fachbereich Physik, der Technischen Universitдt Darmstadt Darmstadt, Germany, 2019.

28. (330) Qianqian Xia, On controlled invariance of regular distributions, Mathematics of Control, Signals, and Systems, volume 33, pp. 79–107, 2021. https://orcid.org/0000-0002-2337-8407, DOI: 10.1007/s00498-020-00275-7. http://arXiv.org e-Print archive, E-print No. arXiv:2111.08850 [math.OC], 2021.

29. (331) Niels Linnemann and James Read, Constructive Axiomatics in Spacetime Physics Part I: Walkthrough to the Ehlers-Pirani-Schild Axiomatisation, An Archive for Preprints in Philosophy of Science, http://philsci-archive.pitt.edu/20056/; http://arXiv.org e-Print archive, E-print No. arXiv:2112.14063 [gr-qc], 2021.

30. (332) Samuel C. Fletcher and James Owen Weatherall, The Local Validity of Special Relativity, Part 1: Geometry, DOI: 10.48550/arXiv.2206.10768, https://doi.org/10.48550/arXiv.2206.10768 http://arXiv.org e-Print archive, E-print No. arXiv:2206.10768v1 [physics.hist-ph], 21 Jun 2022.

31. (333) Henrique Gomes, "Is spacetime locally flat?": a note, Preprint, PhilSci Archive, http://philsci-archive.pitt.edu/21311/, 25 Octiber 2022,

32. (334) Niels Linneman, James Read and Nicholas Teh, The local validity of special relativity: the standard story articulated and defended, DOI: https://doi.org/10.48550/arXiv.2305.0153 http://arXiv.org e-Print archive, E-print No. arXiv:2305.01534v1 [physics.hist-ph], 2 and 15 May 2023.

33. (335) Nathaniel Craig, Yu-Tse Lee, Xiaochuan Lu and Dave Sutherland, Effective Field Theories as Lagrange Spaces, DOI: https://doi.org/10.48550/arXiv.2305.09722 http://arXiv.org e-Print archive, E-print No. arXiv:2305.09722 [hep-th], 16 May 2023.

34. (336) Fabio Sozio and Arash Yavari, A Geometric Field Theory of Dislocation Mechanics, Journal of Nonlinear Science, vol. 33, No. 5, 2023, DOI: 10.1007/s00332-023-09919-9 (http://dx.doi.org/10.1007/s00332-023-09919-9); DOI: https://doi.org/10.48550/arXiv.2307.00164 http://arXiv.org e-Print archive, E-print No. arXiv:2307.00164 [cond-mat.mtrl-sci], 30 June 2023.

35. (337) Поташов И. М., Обобщеный алгоритм вычисление коеффициентов ковариантыйх рядов Тейлора (Generalized algorithm of computation of coefficients of the covariant Taylor series), Вестник Тверского тосударственого университета (Вестник ТеГУ), Серия: Прикладная математика, Вычислителная математика, Выпуск 2, страницы 51–66, Тверь, 2023, DOI: https://doi.org/10.26456/vtpmk670 (Mi vtpmk670); УДК: 514.743, 519.165 URL https://www.mathnet.ru/php/archive.phtml?wshow=paper& jrnid=vtpmk& paperid=670& option_lang=rus Цитат 7.

36. (338) Samuel C. Fletcher, and James Owen Weatherall, The Local Validity of Special Relativity, Part 1: Geometry, Philosophy of Physics 1(1): 7, 1–21, 2023, DOI: https://doi.org/10.31389/pop.6 ; URL: https://philosophyofphysics.lse.ac.uk/articles/10.31389/pop.6

The work [40] is cited by:

1. (339) Marielle Hui Yi Ong, Donaldson’s proof of the Narasimhan-Seshadri Theorem, A thesis submitted for the degree of Master of Philosophy at The University of Queensland in 2018, School of Mathematics and Physics, The University of Quinsland, Austalia, 2018.

The work [41] is cited by:

1. (340) Tunahan Turhan and Nihat Ayyildiz, Characterization of integral curves of a linear vector field in Lorentz 3-space, International Journal of Geometric Methods in Modern Physics, Volume 13, Issue 6, id. 1650073-35, DOI: 10.1142/S0219887816500730, 2016.

The work [42] is cited by:

1. (341) Article in the Wikipedia (French edition, https://fi.wikipedia.org), Gamma-matriisit,

https://fi.wikipedia.org/wiki/Gamma-matriisit

The work [43] is cited by:

1. (342) T. Ryu, T. Tanaka and R. Perna, Population III X-Ray Binaries, Frontier Research in Astrophysics II, held 23-28 May, 2016 in Mondello (Palermo), Italy (FRAPWS2016), 2016. Online at https://pos.sissa.it/cgi-bin/reader/conf.cgi?confid=269, id.40

The work [44] is cited by:

1. (343) Crasmareanu, M. and Frigioiu, C., Unitary vector fields are Fermi-Walker transported along Rytov-Legendre curves, International Journal of Geometric Methods in Modern Physics, Volume 12, Issue 10, 1 November 2015, Article number 1550111, DOI: 10.1142/S021988781550111X.

The work [45] is cited by:

1. (344) Daniel Grumiller (Author, Editor), Anton Rebhan (Editor) and Dimitri Vassilevich (Editor), Fundamental Interactions: A Memorial Volume for Wolfgang Kummer, World Scientific. 2009.

2. (345) Article in the Wikipedia (French edition, https://fi.wikipedia.org), Jauge de Lorenz,

https://fr.wikipedia.org/wiki/Jauge_de_Lorenz,

see also

https://fr.wikipedia.org/wiki/Discussion:Jauge_de_Lorenz

3. (346) Article in http://dictionnaire.sensagent.leparisien.fr Jauge de Lorenz,

http://dictionnaire.sensagent.leparisien.fr/JAUGE DE LORENZ/fr-fr/

4. (347) D. N. Blaschke et. al., Gage fixing in Yang-MIlls Theories, in Fundamental Interactions: A Memorial Volume for Wolfgang Kummer Edited by Daniel Grumiller, Anton Rebhan, and Dimitri Vassilevich, World Scientific, 2009; ISBN: 978-981-4273-07-7 (hardcover), ISBN: 978-981-4468-01-5 (ebook), DOI: http://dx.doi.org/10.1142/9789814277839_fmatter

5. (348) Alessio Rocci, History of Quantum Gravity: from the birth of General Relativity to the end of WWII (1915-1945) (History of Quantum Gravity: from the birth of General Relativity to the end of WWII (1915-1945)), URL: http://paduaresearch.cab.unipd.it/8916/, Note: Presented on 2016-07-30, Universitа degli Studi di Padova, Dipartimento di Fisica e Astronomia ”Galileo Galilei” Supervisor: Kurt Lechner, Ph.D. thesis, 2015

6. (349) Jonas Frafjord, The Vacuum Polarisation Contribution to the Lamb Shift Using Non-Relativistic Quantum Electrodynamics, Ph.D. Thesis, Norwegian University of Science and Technology, Department of Physics, Trondheim, Norway, 2016.

The work [46] is cited by:

1. (350) Tariq Ahmad Shah, Sumeer Gul, and Ramesh C Gaur, Authors self-citation behaviour in the field of Library and Information Science, Aslib Journal of Information Management, Volume 67, Issue 4, pp.458 - 468, 2015; DOI: http://dx.doi.org/10.1108/AJIM-10-2014-0134

2. (351) M. G. Kreines and E. M. Kreines, The control model for the selection of reference collections providing the impartial assessment of the quality of scientific and technological publications by using bibliometric and scientometric indicators, Journal of Computer and Systems Sciences International, Volume 55, Issue 5, pp. 750–766, September 2016.

3. (352) Е. П. Крейнес М. Г. Крейнес, Модель управления выбором референтных коллекций для объективной оценки качества научно-технических публикаций по библиометрическим и наукометрическим показателям, Известий Российской Академии Наук, Теория системый управлений, издательство Российско Академиий Наук, Москва, 2016. ISSN: 0002-3388, DOI: 10.7868/S0002338816040090, УДК: 519.847

The work [47] is cited by:

1. (353) Lorna Wildgaard, A critical cluster analysis of 44 indicators of author-level performance, Journal of Informetrics, Volume 10, Issue 4, Pages 1055-1078, November 2016; DOI: https://doi.org/10.1016/j.joi.2016.09.003

2. (354) Zahid Ashraf Wani and Tazeem Zainab, A review of eminence of scientometric indicators in scientific research productivity, COLLNET Journal of Scientometrics and Information Management, Pages 1-14, Published online: 26 Oct 2017; DOI: http://dx.doi.org/10.1080/09737766.2017.1306179 3. (355) Zahid Ashraf Wani and Tazeem Zainab, Scholarly Content and Its Evolution by

Scientometric Indicators: Emerging Research and Opportunities, IGI Global, 1 edition, 3 August 2018, 201 pages, ISBN13: 9781522559450, ISBN10: 1522559450, EISBN13: 9781522559467, DOI: 10.4018/978-1-5225-5945-0

List of References

[1] Bozhidar Z. Iliev and S. Manov. Deviation equations in spaces with torsion. In Proceedings of the 5-th Soviet (USSR) Gravity Conference “Modern theoretical and experimental problems of relativity theory and gravitation”, page 122, Moscow, 1981. Moscow State University. In Russian. http://arXiv.org e-Print archive, E-print No. gr-qc/0507002 , July 1, 2005. In English.

[2] Bozhidar Z. Iliev and S. Manov. Deviation equations in spaces with affine connection. JINR Communication P2-83-897, Dubna, 1983. In Russian, 14 pp. http://arXiv.org e-Print archive, E-print No. math.DG/0512008, December 1, 2005. In English.

[3] Bozhidar Z. Iliev. On some generalizations of the Jacobi identity. I. Bulletin de la Société des Sciences et des Lettres de Łodz, vol.42, No.21, Série: Recherches sur les Déformations, 14(131):5–11, 1992.

[4] Bozhidar Z. Iliev. Special bases for derivations of tensor algebras. I. Cases in a neighborhood and at a point. JINR Communication E5-92-507, Dubna, 1992. 19 pp. http://arXiv.org e-Print archive, E-print No. math.DG/0303373, March 2003. See also [16].

[5] Bozhidar Z. Iliev. Special bases for derivations of tensor algebras. II. Case along paths. JINR Communication E5-92-508, Dubna, 1992. 16 pp. http://arXiv.org e-Print archive, E-print No. math.DG/0304157, April 2003. See also [19].

[6] Bozhidar Z. Iliev. Special bases for derivations of tensor algebras. III. Case along smooth maps with separable points of selfintersection. JINR Communication E5-92-543, Dubna, 1992. 15 pp. http://arXiv.org e-Print archive, E-print No. math.DG/0305061, May 2003. See also [21].

[7] Bozhidar Z. Iliev. Linear transports along paths in vector bundles. I. General theory. JINR Communication E5-93-239, Dubna, 1993. 22 pp. http://arXiv.org e-Print archive, E-print No. math.DG/0411023, November 1, 2004.

[8] Bozhidar Z. Iliev. Linear transports along paths in vector bundles. II. Some applications. JINR Communication E5-93-260, Dubna, 1993. 24 pp. http://arXiv.org e-Print archive, E-print No. math.DG/0412010, December 1, 2004.

[9] Bozhidar Z. Iliev. Linear transports along paths in vector bundles. III. Curvature and torsion. JINR Communication E5-93-261, Dubna, 1993. 12 pp. http://arXiv.org e-Print archive, E-print No. math.DG/0502008, February 1, 2005.

[10] Bozhidar Z. Iliev. Transports along paths in fibre bundles. General theory. JINR Communication E5-93-299, Dubna, 1993. 26 pp. http://arXiv.org e-Print archive, E-print No. math.DG/0503005, March 1, 2005.

[11] Bozhidar Z. Iliev. Linear transports along paths in vector bundles. IV. Consistency with bundle metrics. JINR Communication E5-94-17, Dubna, 1994. 22 pp. http://arXiv.org e-Print archive, E-print No. math.DG/0505008, May 1, 2005.

[12] Bozhidar Z. Iliev. Transports along paths in fibre bundles. II. Ties with the theory of connections and parallel transports. JINR Communication E5-94-16, Dubna, 1994. 24 pp. http://arXiv.org e-Print archive, E-print No. math.DG/0503006, March 1, 2005.

[13] Bozhidar Z. Iliev. Transports along paths in fibre bundles. III. Consistency with bundle morphisms. JINR Communication E5-94-41, Dubna, 1994. 19 pp. http://arXiv.org e-Print archive, E-print No. dg-ga/9704004, April 1997.

[14] Bozhidar Z. Iliev. Deviation equations in spaces with a transport along paths. JINR Communication E5-94-40, Dubna, 1994. 22 pp. http://arXiv.org e-Print archive, E-print No. math-ph/0303002, March 2003.

[15] Bozhidar Z. Iliev. Centre of mass in spaces with torsion free flat linear connection. JINR Communication E5-95-159, Dubna, 1995. 15 pp. http://arXiv.org e-Print archive, E-print No. gr-qc/0404002, April 2004.

[16] Bozhidar Z. Iliev. Normal frames and the validity of the equivalence principle: I. Cases in a neighborhood and at a point. Journal of Physics A: Mathematical and General, 29(21):6895–6901, 1996. http://arXiv.org e-Print archive, E-print No. gr-qc/9608019, August 1998.

[17] Bozhidar Z. Iliev. Linear transports along paths in vector bundles. V. Properties of curvature and torsion. JINR Communication E5-97-1, Dubna, 1997. 11 pp. http://arXiv.org e-Print archive, E-print No. dg-ga/9709017, September 1997.

[18] Bozhidar Z. Iliev. Transports along maps in fibre bundles. JINR Communication E5-97-2, Dubna, 1997. 19 pp. http://arXiv.org e-Print archive, E-print No. dg-ga/9709016, September 1997.

[19] Bozhidar Z. Iliev. Normal frames and the validity of the equivalence principle: II. The case along paths. Journal of Physics A: Mathematical and General, 30(12):4327– 4336, 1997. http://arXiv.org e-Print archive, E-print No. gr-qc/9709053, September 1997.

[20] Bozhidar Z. Iliev. Is the principle of equivalence a principle? Journal of Geometry and Physics, 24(3):209–222, 1998. http://arXiv.org e-Print archive, E-print No. gr-qc/9806062, June 1998.

[21] Bozhidar Z. Iliev. Normal frames and the validity of the equivalence principle: III. The case along smooth maps with separable points of self-intersection. Journal of Physics A: Mathematical and General, 31(4):1287–1296, January 1998. http://arXiv.org e-Print archive, E-print No. gr-qc/9805088, May 1998.

[22] Bozhidar Z. Iliev. Quantum mechanics from a geometric-observer’s viewpoint. Journal of Physics A: Mathematical and General, 31(4):1297–1305, January 1998. http://arXiv.org e-Print archive, E-print No. quant-ph/9803083, March 1998.

[23] Bozhidar Z. Iliev. On a connection between Lorentzian and Euclidean metrics. Journal of Geometry and Physics, 34(4):321–335, 2000.

[24] Bozhidar Z. Iliev. Fibre bundle formulation of nonrelativistic quantum mechanics. I. Introduction. The evolution transport. Journal of Physics A: Mathematical and General, 34(23):4887–4918, 2001.

[25] Bozhidar Z. Iliev. Fibre bundle formulation of nonrelativistic quantum mechanics. II. Equations of motion and observables. Journal of Physics A: Mathematical and General, 34(23):4919–4934, 2001.

[26] Bozhidar Z. Iliev. Fibre bundle formulation of nonrelativistic quantum mechanics. III. Pictures and integrals of motion. Journal of Physics A: Mathematical and General, 34(23):4935–4950, 2001.

[27] Bozhidar Z. Iliev. Fibre bundle formulation of nonrelativistic quantum mechanics. IV. Mixed states and evolution transport’s curvature. International Journal of Modern Physics A, 17(2):229–243, 2002. http://arXiv.org e-Print archive, E-print No. quant-ph/9901039, January 1999.

[28] Bozhidar Z. Iliev. Fibre bundle formulation of nonrelativistic quantum mechanics. V. Interpretation, summary, and discussion. International Journal of Modern Physics A, 17(2):245–258, 2002. http://arXiv.org e-Print archive, E-print No. quant-ph/9902068, February 1999.

[29] Bozhidar Z. Iliev. On metric-connection compatibility and the signature change of space-time. Physica Scripta, 66(6):401–409, 2002. http://arXiv.org e-Print archive, E-print No. gr-qc/9802058, February 1998.

[30] Bozhidar Z. Iliev. Fibre bundle formulation of relativistic quantum mechanics. http://arXiv.org e-Print archive, E-print No. quant-ph/0201085, January 2002.

[31] Bozhidar Z. Iliev. On operator differentiation in the action principle in quantum field theory. In Stancho Dimiev and Kouei Sekigava, editors, Proceedings of the 6th International Workshop on Complex Structures and Vector Fields, 3–6 September 2002, St. Knstantin resort (near Varna), Bulgaria, “Trends in Complex Analysis, Differential Geometry and Mathematical Physics”, pages 76–107. World Scientific, New Jersey-London-Singapore-Hong Kong, 2003. http://arXiv.org e-Print archive, E-print No. hep-th/0204003, April 2002.

[32] Bozhidar Z. Iliev. Normal frames for derivations and linear connections and the equivalence principle. Journal of Geometry and Physics, 45(1–2):24–53, February 2003. http://arXiv.org e-Print archive, E-print No. hep-th/0110194, October 2001.

[33] Bozhidar Z. Iliev. Equivalence principle in classical electrodynamics. In Andrew Chubykalo, Vladimir Onoochin, Augusto Espinoza, and Roman Smirnov-Rueda, editors, Has the last word been said on classical electrodynamics? – New Horizons, pages 385–403. Rinton Press, Princeton, NJ, 2004. ISBN-13: 978-1589490369, ISBN-10: 1589490363 http://arXiv.org e-Print archive, E-print No. gr-qc/0020303002, March 2003.

[34] Bozhidar Z. Iliev. Pictures and equations of motion in Lagrangian quantum field theory. In Charles V. Benton, editor, Studies in Mathematical Physics Research, pages 83–125. Nova Science Publishers, Inc., New York, 2004. 1594540276, ISBN-13: 978-1594540271 http://arXiv.org e-Print archive, E-print No. hep-th/0302002, February 2003.

[35] Bozhidar Z. Iliev. Lagrangian quantum field theory in momentum picture. II. Free spinor fields. http://arXiv.org e-Print archive, E-print No. hep-th/0405008, May 1, 2004.

[36] Bozhidar Z. Iliev. Normal frames and linear transports along paths in vector bundles. JP Journal of Geometry and Topology, 5(3):187–250, 2005. http://arXiv.org e-Print archive, E-print No. gr-qc/9809084, September 1998 (last revision: March 2005).

[37] Bozhidar Z. Iliev. Links between connections, parallel transports, and transports along paths in differentiable fibre bundles. International Journal of Geometric Methods in Modern Physics, 2(5):823–838, 2005. Talk at the International Workshop on “Advanced Geometric Methods in Physics”, Florence, Italy, 14 – 18 April, 2005. http://arXiv.org e-Print archive, E-print No. math.DG/0504010, April 1, 2005.

[38] Bozhidar Z. Iliev. Lagrangian quantum field theory in momentum picture. III. Free vector fields. http://arXiv.org e-Print archive, E-print No. hep-th/0505007, May 1, 2005.

[39] Bozhidar Z. Iliev. Handbook of Normal Frames and Coordinates, volume 42 of Progress in Mathematical Physics. Birkhäuser, Basel-Boston-Berlin, 2006. 468 pages, Hardcover, ISBN (ISBN-10): 3-7643-7618-X, ISBN-13: 978-3-7643-7618-5, EAN: 9783764376185; http://arXiv.org e-Print archive, E-print No. math.DG/0610037, October 1, 2006.

[40] Bozhidar Z. Iliev. On the axiomatization of “parallel transport” and its generalization. Tensor, N. S., 62(2):170–181, 2006. Report presented at the 8th International Conference of Tensor Society on “Differential geometry, functional and complex analysis, informatics and their applications”, Varna, Bulgaria, 22 – 26 August, 2005 http://arXiv.org e-Print archive, E-print No. math-ph/0510005, October 1, 2005.

[41] Bozhidar Z. Iliev and Rahula M. Linear vector fields and exponential law. http://arXiv.org e-Print archive, E-print No. math-ph/0604005, March 1, 2006.

[42] Bozhidar Z. Iliev. Langrangian Quantum Field theory in Momentum Picture. Free fields. Nova Science Publishers, New York, 2008. 306 pages, Hardcover, ISBN-13: 978-1-60456-170-8, ISBN-10: 1-60456-170-X.

[43] Bozhidar Z. Iliev. Momentum picture of motion in Lagrangian quantum field theory. International Journal of Theoretical Physics, Group Theory, and Nonlinear Optics, 12(2):103–121, 2008. http://arXiv.org e-Print archive, E-print No. hep-th/0311003, November, 2003.

[44] Bozhidar Z. Iliev. Connections, parallel transports and transports along paths. International Journal of Geometric Methods in Modern Physics, 5(7):1137–1162, 2008. http://arXiv.org e-Print archive, E-print No. math.DG/0504010v4, May 1, 2008.

[45] Bozhidar Z. Iliev. The “Lorenz gauge” is named in honour of Ludowig Valentin Lorenz! http://arXiv.org e-Print archive, E-print No. arXiv:0803.0047 [physics.hist-ph], May 1, 2008.

[46] Bozhidar Z. Iliev. Measuring the evaluation and impact of scientific works and their authors. http://arXiv.org e-Print archive, E-print No. arXiv:1311.6948 [cs.DL], November 27, 2013.

[47] Bozhidar Z. Iliev. Modern bibliometric indicators and achievements of authors. Journal of Geometry and Symmetry in Physics, 33:113–128, 2014.

Produced on 19 November 2023 at 11:33.